### v1

# Generalizable Deepfake Detection via Effective Local-Global Feature Extraction

Abstract
The rapid advancement of GANs and diffusion models has led to the generation of increasingly realistic fake images, posing significant hidden dangers and threats to society. Consequently, deepfake detection has become a pressing issue in today’s world. While some existing methods focus on forgery features from either a local or global perspective, they often overlook the complementary nature of these features. Other approaches attempt to incorporate both local and global features but rely on simplistic strategies, such as cropping, which fail to capture the intricate relationships between local features. To address these limitations, we propose a novel method that effectively combines local spatial-frequency domain features with global frequency domain information, capturing detailed and holistic forgery traces. Specifically, our method uses Discrete Wavelet Transform (DWT) and sliding windows to tile forged features and leverages attention mechanisms to extract local spatial-frequency domain information. Simultaneously, the phase component of the Fast Fourier Transform (FFT) is integrated with attention mechanisms to extract global frequency domain information, complementing the local features and ensuring the integrity of forgery detection. Comprehensive evaluations on open-world datasets generated by 34 distinct generative models demonstrate a significant improvement of 2.9% over existing state-of-the-art methods.
1 Introduction
With the rapid advancement of deep learning, generative models such as GANs and diffusion models have revolutionized content creation, finding widespread applications across industries including entertainment and media. However, this progress has also introduced significant challenges. The increasing sophistication of deepfake technology has allowed the generation of highly realistic fake images and videos, making them increasingly difficult to detect and posing serious risks to societal trust, individual privacy, and security.
In response, various deepfake detection mechanisms have been conceived. Early methods primarily focused on analyzing image features, such as texture, color, and lighting Yu et al. (2020); Liu et al. (2020), to identify forgery traces. However, these approaches were often limited to training and testing within a single domain, overlooking the critical issue of cross-domain generalization. In real-world scenarios, it is essential for detection models to effectively generalize to unseen deepfake sources, including those involving unfamiliar generative models or novel categories.
Figure 1: Comparison with Previous Methods. Previous methods often extract only local information or global information to detect artifact information. However, these methods are not enough to fully capture effective artifact information. Recent methods also combine local and global information to detect forged images, but they only extract local information through simple cropping, destroying the relationship between local information. Compared with them, our method uses DWT and sliding window tiling to reasonably and effectively extract local time-frequency information, and combines the phase part of FFT to extract global frequency domain information.
In recent years, substantial efforts have been devoted to forgery detection, particularly to identifying deepfake sources that are challenging to detect in real-world scenarios, as illustrated in Figure 1 (a). To enhance generalization performance, some studies emphasize fine-grained features, such as local artifacts Haliassos et al. (2021); Chen et al. (2022) and the relationships between pixel points Chen et al. (2021); Tan et al. (2024b). Others focus on global artifact traces, including global spatial information Tan et al. (2023); Ojha et al. (2023) and frequency-domain features Tan et al. (2024a). However, prioritizing either local or global information in isolation may result in the loss of complementary details, thereby limiting generalization capabilities. To address this limitation, some methods integrate both local and global features. For example, Chen and Yang (2020) extract local features through clipping in the spatial domain while leveraging global features in the frequency space. Similarly, Wu et al. (2024) employ a local attention module and facilitates interaction with features from distant adjacent PPG graphs to capture local and global information, respectively. However, these approaches often rely on coarse image block segmentation for local feature extraction, which can lead to forced separation and loss of relevant feature relationships.
Accordingly, we propose a simple yet efficient feature extractor to capture global frequency domain and local spatial-frequency domain information, as illustrated in Figure 1 (b), which is plug-and-play and can be integrated with any CNN classifier. Specifically, our method introduces two key modules, DWT-Based Local Spatial-Frequency Domain Feature Capture Block (LoSFB) and FFT-Based Global Frequency Domain Feature Capture Block (GloFB), both using the Window Attention Block as the feature extractor. After performing the DWT, we concatenate features from four different frequency bands, each with dimensions $h\times w$. Then a $2\times 2$ sliding window is applied to extract the tiled features from each frequency band. Finally, the features within the window are tiled again, resulting in new features with dimensions $4\times(h\times w)$. At this time, when we extract the $4\times 4$ window features, they not only obtain features in four frequency domains, but also retain the local spatial-frequency information within the window, greatly improving the effect of artifact detection. Additionally, we incorporate the phase information from the FFT as a complement to the DWT, enabling the extraction of global frequency domain information. This helps guide the network to effectively learn global artifact patterns and identify more comprehensive traces of forgery. To fully assess the generalizability of our method, we evaluate it on large image datasets generated by 34 distinct models [^1] ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, SAN, AttGAN, BEGAN, CramerGAN, InfoMaxGAN, MMDGAN, RelGAN, S3GAN, SNGAN, STGAN, ADM, DDPM, IDDPM, LDM, PNDM, VQDiffusion, Stable-Diffusion v1, Stable-Diffusion v2, DALLE, Glide, Midjourney, Stable-Diffusion v1.4, Stable-Diffusion v1.5, Wukong, DALLE-2. Comprehensive experiments demonstrate that our method generalizes well across diverse and unseen sources, demonstrating a significant improvement over state-of-the-art methods by 2.9%. Our paper makes the following contributions:
  • We propose a plug-and-play feature extractor that seamlessly integrates local spatial-frequency and global frequency-domain information. This method can be effortlessly incorporated into any CNN classifier and effectively improves the model’s generalization capability.
  • We use sliding window interpretability to extract local spatial frequency information, and creatively use only phase spectrum to obtain global features to capture broader forgery artifacts.
  • Our experiments validate the effectiveness of the proposed method, which exhibits strong generalization across image detection tasks involving 34 different generation models. The method achieves a notable 2.9% improvement over state-of-the-art approaches, highlighting its superiority and robustness.
2 Related Work
In this section, we present a concise review of deepfake detection methods, mainly dividing them into three main groups: deepfake detection based on local information, deepfake detection based on global information, and deepfake detection based on local and global information fusion.
2.1 Deepfake Detection Based on Local Information
Early researches on deepfake detection often focus on specific facial regions, such as the eyes and lips, to identify forgery traces Haliassos et al. (2021); Chen et al. (2022). However, as forgery techniques have advanced, these simple traces have been increasingly refined and obscured, rendering biometric-based detection methods insufficient for identifying sophisticated forged images. Experiments have shown that forgery traces are often embedded in fine-grained details, prompting a shift in focus toward extracting local artifact information. For instance, Chai et al. (2020) use limited receptive fields to capture local features, while Zhao et al. (2021) employ multiple spatial attention heads to enhance the network’s focus on distinct local regions. Other studies analyze forgery traces by examining relationships between local pixels. Chen et al. (2021) propose learning the similarity of local regions for a fine-grained assessment of patch similarity across scales, and Tan et al. (2024b) improve forgery detection generalization by identifying upsampling artifacts in nearby pixel grids. While capturing local artifact details, these methods overlook the fact that the overall image may still contain forged features, which are essential for generalization.
2.2 Deepfake Detection Based on Global Information
Early studies leveraged spatial information from generated images to detect fake content. For example, Yu et al. (2020); Liu et al. (2020) distinguish real and fake images using global information, such as color space and global texture. Some studies Yu et al. (2019) use GANs to generate global fingerprint information that may assist in detection. Additionally, Tan et al. (2023) and Ojha et al. (2023) extract artifact representations using global feature maps and global feature gradients, respectively. Moreover, since GAN-generated images often rely on upsampling operations that introduce artifacts at the frequency domain level, several studies have explored forgery detection in the frequency domain. For instance, Tan et al. (2024a) use FFT to extract global high-frequency information, uncovering more refined forgery features. While these methods capture sufficient global artifact information, they overlook the fact that effective artifact features often reside in subtle, localized regions. The neglect of local artifact information can significantly impair the generalization capabilities of forgery detection.
2.3 Deepfake Detection Based on Local and Global Information Fusion
Whether relying on global or local information alone, these methods detect forgery using only a single category of features, overlooking the significant information gain that arises from combining complementary features at different levels. Chen and Yang (2020) capture local features through spatial domain cropping and global features in the frequency domain. Xue et al. (2023) employ separate branches for local area feature extraction and global feature extraction, using the latter to obtain residual features. Wu et al. (2024) introduce a local attention module that interacts with distant neighboring PPG graphs to gather both local and global information. While these studies incorporate both global and local information, their approach to processing local information is limited by artificial segmentation of the face into regions, which leads to the loss of important forged details.
3 Main Backbone
The overall structure of our method is shown in Figure 2. Inspired by Mao et al. (2024), we propose an efficient local-global feature extraction framework designed to capture both global and local information. Specifically, the framework consists of two main branches: LoSFB and GloFB. Both branches utilize the Window Attention Block to extract the most relevant features, which are then fused and passed to the classifier for final classification. Notably, our approach is compatible with any classifier, allowing us to focus primarily on the feature extraction process.
Figure 2: Architecture of Our Method for Generalizable Deepfake Detection. In order to capture more comprehensive and reasonable forgery features, we use LoSFB and GloFB to capture local and global information, respectively. Specifically, we use DWT and sliding window tiling to construct local spatial-frequency features, and only use the phase part of FFT to construct global frequency domain features. Both of them use Window Attention Block to extract features. In addition, in order to weaken the information weight of global features, we add hyperparameters $\lambda$ to the global features when merging the two types of features. The final extracted features are passed into a classifier.
3.1 Window Attention Block
The Window Attention Block (WAttention) consists of two main components: (i) the Local Attention Block (LoAttention) and (ii) the Global Window MLP Layer (GMLP). Before feeding the features into LoAttention and GMLP, we preprocess them to enhance their representational quality. A $1\times 1$ convolution layer is applied to extract cross-channel features, following by a $3\times 3$ depthwise separable convolution to better capture the local features within each channel. The resulting convolved features are denoted as $X_{in}\in{\mathbb{R}}^{B\times C\times H\times W}$.
LoAttention. Unlike traditional attention mechanisms, which are typically designed to capture global information, we employ a sliding window approach that divides the features into windows of size $b\times b$. By restricting the attention module to these local windows, the focus shifts to effectively extracting local features. The features are then divided into $Q,K,V\in{\mathbb{R}}^{S\times C\times N}$, where $S=B\times\frac{H}{b}\times\frac{W}{b}$ and $N=b\times b$, as required by the attention module. Since the attention operates locally within each window, the features within window $i$ are represented as $Q_{i},K_{i},V_{i}\in{\mathbb{R}}^{C\times N}$. These localized features are fed into the attention module for computation:
$A_{i}=\text{Attention}(Q_{i},K_{i},V_{i})=\text{SoftMax}(Q_{i}K_{i}^{\top})V_{i}.$ (1)
The results of all window attention operations can be expressed as $A=\{A_{1},A_{2},\cdots,A_{S}\}\in{\mathbb{R}}^{S\times C\times N}$.
GMLP. To preserve global correlations while capturing fine-grained features, we integrate a global MLP within each window. This aims to address the limitations of local windows in modeling global relationships, ensuring a seamless fusion of local details and global context. Specifically, we reuse the value $V_{i}$ from LoAttention and process it through the global MLP, applying a nonlinear transformation to enhance its representational capacity. The result $\hat{V_{i}}$ is formally expressed as:
$\text{GMLP}(V_{i})=G(\text{Linear}(V_{i})),$ (2)
where $G(\cdot)$ denotes the GELU function. The GMLP recombines and redistributes the local window information, enabling the mixing of local frequency domain features and enhancing the learning of global features.
After combining the outputs from LoAttention and GMLP via a dot product, we use the inverse window operation to restore the feature to its original size $B\times C\times H\times W$, followed by a convolution to map the features back to the original feature dimensions. The final output of the WAttention can be expressed as:
$\text{X}_{W}=\text{conv}(\sum_{i=1}^{s}\text{Attention}(Q_{i},K_{i},V_{i})\cdot\sum_{i=1}^{s}G(\text{Linear}(V_{i}))).$ (3)
3.2 DWT-Based Local Spatial-Frequency Domain Feature Capture Block
Previous methods often leverage the Fast Fourier Transform (FFT) or Discrete Cosine Transform (DCT) to capture relationships between adjacent nodes in the frequency domain. However, these approaches can lead to the loss of positional information, which is crucial for extracting detailed features. To overcome this limitation, we utilize the Discrete Wavelet Transform (DWT), which decomposes the image into the spatial-frequency domain, represented as:
$\text{DWT}(I)=\{I_{LL},I_{LH},I_{HL},I_{HH}\},$ (4)
where $I_{LL},I_{LH},I_{HL},I_{HH}\in{\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}}$ represent the low-frequency component of the input and the high-frequency components in the vertical, horizontal, and diagonal directions, respectively. To obtain more comprehensive information, we fuse the data from the four frequency domains, resulting in new input features $\hat{I}\in{\mathbb{R}^{4\times C\times\frac{H}{2}\times\frac{W}{2}}}$. As illustrated in Figure 3, to extract detailed features, we divide each channel’s features into $2\times 2$ windows, with each window capturing localized spatial-frequency information within its respective frequency band. Each window is treated as a small square and tiled into rows within each band. This arrangement ensures that features within the $2\times 2$ windows are preserved in a structured format. Then we tile the features in the window here by rows. At this time, the features from each frequency band are tiled into one row in a specific manner, such that the four adjacent feature values correspond to the $2\times 2$ window from the original image. Additionally, as shown in Figure 3, we integrate features from all frequency bands to generate new input features ${\hat{I}}_{\text{input}}\in{\mathbb{R}}^{C\times 4\times\frac{H\times W}{4}}$. To further enhance the network’s performance, we apply LayerNorm Lei Ba et al. (2016), which ensures that the input data remains within a stable range, reduces dependency on initialization, and improves the stability of model training.
The feature ${\hat{I}}_{\text{input}}$ is then fed into WAttention with a sliding window size set to $b=4$. As depicted in Figure 3, within each $4\times 4$ window, the representation includes not only the four adjacent feature values from the original image but also the feature information across four distinct frequency bands. This configuration provides WAttention with comprehensive and fine-grained spatial-frequency domain information, enabling more detailed feature extraction. Following this process, we apply the IDWT to reconstruct the features back into spatial domain representations, denoted as ${\hat{I}_{\text{DWT\_output}}}$.
3.3 FFT-Based Global Frequency Domain Feature Capture Block
Although artifact information is often concentrated in finer details, forged features can also be embedded in the global information of an image, remaining imperceptible to the naked eye. While DWT and sliding window tiling are effective for extracting local features, we also use FFT capture global frequency domain information from a complementary perspective. In this module, FFT is employed to transform the image into the frequency domain at the pixel level, enabling the extraction of global frequency domain features. Inspired by Zhao et al. (2024), we recognize that amplitude primarily reflects properties such as color and brightness, while phase encodes finer details, including potential artifacts. Based on this insight, we focus on extracting the phase features as input to WAttention, setting the window size to $b=8$. The amplitude remains unmodified during this process. After training the phase features, we use IFFT to recombine them with the original amplitude, transforming the features back into the spatial domain. The resulting features are denoted as ${\hat{I}_{\text{FFT\_output}}}$.
Finally, we set a hyperparameter $\lambda$ to minimize the impact of global information extraction. The final feature is:
${\hat{X}_{\text{output}}}=(1-\lambda)*\hat{I}_{\text{DWT\_output}}+\lambda*\hat{I}_{\text{FFT\_output}}.$ (5)
Figure 3: After DWT, sliding window tiling is used to tile the features of each channel into a shape of $4\times(h\times w)$. In the subsequent LoAttention module, $4\times 4$ features are extracted as shown in the blue box, which not only contains local spatial frequency information, but also contains all different frequency band features.
4 Experiments
To verify the excellent deepfake detection performance of our method, we followe the paradigms of Ojha et al. (2023) and Wang et al. (2020).
4.1 Dataset
Training set. To ensure a consistent basis for comparison, we use the training set of ForenSynthsWang et al. (2020). The training set consists of 20 different categories, each containing 18,000 synthetic images generated using ProGAN, and an equal number of real images from the LSUN dataset. Following the paradigms of Ojha et al. (2023), we adopt specific 4-class training settings, denoted as (car, cat, chair, horse).
Testing set. In order to comprehensively compare the effectiveness of our proposed method, we select different GAN and diffusion datasets for testing, which contain a wide range of real-world images.
  • ForenSynths: The test set comprises 10 subsets, including: ProGAN Karras (2017), StyleGAN Karras (2019), StyleGAN2 Karras et al. (2020), BigGAN Brock (2018), CycleGAN Zhu et al. (2017), StarGAN Choi et al. (2018), GauGAN Park et al. (2019), Deepfake Rossler et al. (2019), WFIR and SAN Dai et al. (2019).
  • GANGen-Detection Tan et al. (2024a): To evaluate more realistic scenarios, we extend our evaluation using 9 additional GAN-generated images. There are 4K test images for each model, with equal numbers of real and fake images.
  • DiffusionForensics Wang et al. (2023): To expand the testing scope, we adopt the diffusions dataset of DIRE Wang et al. (2023) for evaluation, including ADM Dhariwal and Nichol (2021), DDPM Ho et al. (2020), IDDPM Nichol and Dhariwal (2021), LDM Rombach et al. (2022), PNDM Liu et al. (2022), VQ-Diffusion Gu et al. (2022), Stable Diffusion v1 Rombach et al. (2022), Stable Diffusion v2 Rombach et al. (2022). The real images are sampled from LSUN Yu et al. (2015) and ImageNet Russakovsky et al. (2015) datasets.
  • UniversalFakeDetect Ojha et al. (2023): This test set contains images generated from Glide Nichol et al. (2021), DALL-Emini Ramesh et al. (2021), LDM Rombach et al. (2022). It adopts images of LAION Schuhmann et al. (2021) and ImageNetRussakovsky et al. (2015) datasets as the real data.
  • AIGCDetectBenchmark Zhu et al. (2024): This dataset mainly adopts diffusion model to generate images, and is generated using more realistic advanced methods, including Midjourney, SDv1.4, SDv1.5 Rombach et al. (2022), ADM Dhariwal and Nichol (2021), GLIDE Nichol et al. (2021), Wukong, VQDM Gu et al. (2022), which is a huge challenge for deepfake detection.
4.2 Implementation Details
We design a lightweight CNN network using convolutional layers and Resnet network as classifiers. The network is trained using the Adam optimizerKingma (2014) with a learning rate of $2\times 10^{-4}$. The batchsize is set to 128, and we train the network for 90 epochs. We also use a learning rate decay strategy that reduces the learning rate by twenty percent after every ten epochs. What’s more, the hyperparameter $\lambda$ is set to 0.2. Our approach is implemented using PyTorch on an NVIDIA RTX A6000 GPU. To evaluate the performance of the proposed method, we follow the evaluation metrics used in baselines Ojha et al. (2023); Tan et al. (2024a), which include average precision score (A.P.) and accuracy (Acc.).
4.3 Quantitative analysis
We compare with the previous methods: CNNDetection Wang et al. (2020), Patchfor Chai et al. (2020), F3Net Qian et al. (2020), FrePGAN Jeong et al. (2022), GANDetection Mandelli et al. (2022), LGrad Tan et al. (2023), Ojha Ojha et al. (2023), FreqNet Tan et al. (2024a) and NPR Tan et al. (2024b). We conduct two and three sets of experiments on GAN and diffusion respectively.
GAN. As shown in Table 1 and 2, our detection on GAN has achieved good results. Table 1 presents a detailed comparison of detector performance on the test set of ForenSynths Wang et al. (2020), highlighting higher mean accuracy (Acc.) and comparable mean average precision (A.P.) metrics. For unavailable test results, we use ”-” as a placeholder and exclude these values from the average calculations. Our method achieves an average accuracy of 86.9%, which is 1% and 1.7% higher than the latest methods, NPR and FreqNet, underscoring its effectiveness in generalizable deepfake detection. Furthermore, to evaluate cross-GAN-source generalization, we conducted tests on the GANGen-Detection dataset Tan et al. (2024a), as summarized in Table 2. Our method achieves an impressive average accuracy of 96.5%, significantly outperforming the best-performing baselines, NPR Tan et al. (2024b) and FreqNet Tan et al. (2024a), which achieved 94.8% and 94.0%, respectively.
Diffusion. To further evaluate the generalization ability of our method, we train on ProGAN and test on images generated by various diffusion models. The results are presented in Tables 3 and 4. Despite being trained exclusively on ProGAN-generated images, our method demonstrates strong generalization capabilities across diverse diffusion models. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 98.5% and 99.9%, respectively. Compared to NPR and FreqNet, our method improves mean Acc. by 2.9% and 17.8%, respectively. Additionally, on the UniversalFakeDetect dataset Ojha et al. (2023), our method achieves an average accuracy of 95.0%, representing a notable improvement over existing approaches.
With the rapid advancements in tools like Midjourney and diffusion models, increasingly realistic fake images are being generated, posing significant challenges for deepfake detection. To address this, we conduct additional evaluations on the AIGCDetectBenchmark dataset Zhu et al. (2024), excluding some subsets of data previously tested. The results, summarized in Table 5, highlight the robustness of our approach. Our method achieves a notable improvement in mean accuracy, outperforming NPR and FreqNet by 3.7% and 7.8%, respectively, and reaching an accuracy of 77.6%.
Method ProGAN GauGAN StyleGAN2 BigGAN StyleGAN CycleGAN StarGAN Deepfake WFIR SAN Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
CNNDetection 91.4 99.4 63.9 92.2 76.4 97.5 52.9 73.3 63.8 91.4 72.7 88.6 63.8 92.2 51.7 62.3 91.7 96.3 48.7 59.5 67.7 85.3
Patchfor 97.8 100.0 57.2 55.4 83.6 98.5 64.7 69.5 82.6 93.1 74.5 87.2 100.0 100.0 85.0 93.2 - - 75.3 76.3 80.1 85.9
F3Net 99.4 100.0 58.1 56.1 88.0 99.8 65.3 69.9 92.6 99.7 76.4 84.3 100.0 100.0 63.5 78.8 - - 47.3 46.7 76.7 81.7
FrePGAN 99.0 99.9 60.3 77.1 84.1 98.6 69.2 71.7 80.7 89.6 71.1 74.4 99.9 100.0 70.9 91.9 - - - - 79.3 87.9
LGrad 99.9 100.0 72.4 79.3 96.0 99.9 82.9 90.7 94.8 99.9 85.3 94.0 99.6 100.0 58.0 67.9 60.5 65.1 50.0 51.4 79.9 84.8
Ojha 100.0 100.0 99.7 100.0 75.7 97.8 95.1 99.3 84.4 97.1 98.7 99.8 99.9 99.4 67.4 82.0 64.2 68.7 56.5 82.2 84.2 92.6
FreqNet 99.6 100.0 93.4 98.6 88.0 99.5 90.5 96.0 90.2 99.7 95.8 99.6 85.7 99.8 88.9 94.4 48.0 49.6 71.9 80.1 85.2 91.4
NPR 99.9 100.0 80.9 83.0 99.6 100.0 84.0 85.6 98.1 99.8 95.2 98.1 99.8 100.0 77.2 76.0 60.6 62.9 64.4 66.0 85.9 87.1
ours 99.9 100.0 86.3 97.4 99.9 100.0 87.1 93.6 99.9 100.0 97.1 98.9 90.8 95.9 64.5 72.2 77.8 87.9 65.6 70.6 86.9 91.7
Table 1: Cross-GAN-Sources Evaluation on the test set of ForenSynths Wang et al. (2020). Bold and underline represent the best and second-best performance, respectively.
Method BEGAN SNGAN CramerGAN MMDGAN RelGAN STGAN S3GAN AttGAN InfoMaxGAN Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
CNNDetection 50.2 44.9 62.7 90.4 81.5 97.5 72.9 94.4 53.3 82.1 63.0 92.7 55.2 66.1 51.1 83.7 71.1 94.7 62.3 82.9
Patchfor 97.1 100.0 97.6 99.8 97.8 99.9 97.9 100.0 99.6 100.0 92.7 99.8 66.8 68.8 68.0 92.9 93.6 98.6 90.1 95.4
F3Net 87.1 97.5 51.6 93.6 89.5 99.8 73.7 99.6 98.8 100.0 60.3 99.9 65.4 70.0 85.2 94.8 67.1 83.1 75.4 93.1
GANDetection 67.9 100.0 66.7 90.6 67.8 99.7 67.7 99.3 60.9 86.2 69.6 97.2 69.6 83.5 57.4 75.1 67.6 92.4 66.1 91.6
LGrad 69.9 89.2 78.0 87.4 50.3 50.3 57.5 67.3 89.1 99.1 54.8 68.0 78.5 86.0 68.6 93.8 71.1 82.0 68.6 80.8
Ojha 88.9 96.3 87.7 96.5 89.8 99.1 89.7 99.0 94.0 98.1 82.6 91.5 94.7 98.9 90.0 96.8 87.4 96.5 89.4 97.0
FreqNet 98.8 100.0 85.4 90.4 95.2 98.2 95.2 98.2 100.0 100.0 98.8 100.0 88.3 94.3 89.8 98.8 94.0 97.5 94.0 97.5
NPR 99.6 99.9 93.3 97.3 98.5 98.3 98.5 98.3 99.8 100.0 99.6 100.0 79.8 78.9 92.5 98.6 91.5 97.2 94.8 96.5
ours 97.7 99.8 96.4 99.6 98.2 99.8 98.8 99.9 98.8 88.8 99.4 100.0 83.4 90.2 98.7 99.8 96.8 99.7 96.5 98.7
Table 2: Cross-GAN-Sources Evaluation on the 9 GANs dataset Tan et al. (2024a).
Method DDPM VQ-Diffusion SDv1 ADM SDv2 IDDPM PNDM LDM Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
CNNDetection 62.7 76.6 50.0 71.0 38.0 76.7 53.9 71.8 52.0 90.3 50.2 82.7 50.8 90.3 50.4 78.7 51.0 90.3
Patchfor 62.3 97.1 100.0 100.0 90.7 99.8 77.5 93.9 94.8 100.0 50.0 91.6 50.2 99.9 99.5 100.0 78.1 97.8
F3Net 84.7 99.4 100.0 100.0 73.4 97.2 80.9 96.9 99.8 100.0 74.7 98.9 72.8 99.5 100.0 100.0 85.8 99.0
GANDetection 62.3 46.4 51.1 51.2 39.8 65.6 51.1 53.1 50.1 36.9 50.2 63.0 50.6 79.0 51.6 48.1 50.8 55.4
LGrad 99.9 100.0 96.2 100.0 90.4 99.4 86.4 97.5 97.1 100.0 66.1 92.8 69.5 98.5 99.7 100.0 88.2 98.5
Ojha 72.9 78.8 77.7 99.2 59.0 93.0 88.7 98.6 53.3 87.6 73.4 97.1 86.2 99.2 50.7 87.9 70.2 92.5
FreqNet 91.4 99.8 100.0 100.0 63.9 98.1 67.2 91.3 81.8 98.4 59.0 97.3 85.2 99.8 98.9 100.0 80.9 98.1
NPR 99.4 100.0 100.0 100.0 97.3 99.8 88.7 98.6 96.4 100.0 87.5 98.3 97.5 100.0 100.0 100.0 95.8 99.6
ours 99.4 100.0 99.9 100.0 95.6 99.7 96.3 99.4 99.2 100.0 99.4 100.0 99.4 100.0 99.8 100.0 98.7 99.9
Table 3: Cross-Diffusion-Sources Evaluation on the test of DiffusionForensics Wang et al. (2023).
Method Guided Glide_50_27 Glide_100_10 Glide_100_27 LDM_100 LDM_200 LDM_200_cfg dalle Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
CNNDetection 54.9 66.6 54.2 76.0 53.3 72.9 53.0 71.3 51.9 63.7 52.0 64.5 51.6 63.1 51.8 61.3 52.8 67.4
Patchfor 74.2 81.4 84.9 98.8 87.3 99.7 82.8 99.1 95.8 99.8 95.6 99.9 94.0 99.8 79.8 99.1 86.8 97.2
F3Net 69.2 70.8 88.5 95.4 88.3 95.4 87.0 94.5 74.1 84.0 73.4 83.3 80.7 89.1 71.6 79.9 79.1 86.5
GANDetection 50.1 51.0 51.7 53.5 51.2 52.6 51.1 51.9 54.7 65.8 54.9 65.9 53.8 58.9 67.2 83.0 54.3 60.1
LGrad 71.8 76.0 91.6 96.0 90.9 95.5 88.5 94.2 95.7 99.3 94.7 99.1 95.1 99.0 89.9 97.8 89.8 94.6
Ojha 69.7 87.6 79.1 99.8 77.9 94.5 78.5 95.3 95.0 99.3 99.0 99.3 74.0 92.5 87.3 97.5 82.6 95.7
FreqNet 67.3 75.7 86.7 96.3 87.9 96.4 84.5 96.0 97.9 99.9 97.5 99.9 97.4 99.9 97.4 99.8 89.6 95.5
NPR 74.8 78.8 98.3 99.8 98.7 99.7 98.2 99.8 98.8 99.8 99.0 99.8 98.8 99.7 95.9 99.7 94.8 97.0
ours 82.4 87.1 98.3 99.7 97.6 99.8 97.6 99.6 99.4 100.0 99.3 100.0 99.0 99.9 86.0 95.6 95.0 97.7
Table 4: Cross-Diffusion-Sources Evaluation on the diffusion test set of Ojha Ojha et al. (2023).
Method ADM Glide Midjourney SDv1.4 SDv1.5 VQDM Wukong DALLE2 Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
LGrad 65.8 70.2 60.6 73.8 64.6 70.3 65.9 68.7 66.3 69.1 67.2 70.6 63.2 65.0 66.9 86.9 65.1 71.8
Ojha 67.6 87.2 61.7 85.5 56.7 74.5 63.1 87.2 62.8 86.3 84.9 96.8 71.2 90.9 49.5 60.2 73.5 83.6
FreqNet 83.3 91.5 81.7 88.9 69.9 79.0 64.3 74.3 64.9 75.5 81.7 89.6 57.7 67.0 55.1 54.6 69.8 77.6
NPR 70.6 75.1 77.2 83.2 76.6 82.6 76.8 82.8 78.0 83.8 75.2 77.7 74.5 78.2 61.9 71.6 73.9 79.4
ours 80.2 83.7 80.8 93.0 73.0 79.4 78.4 84.1 77.8 87.9 83.2 88.7 75.5 78.0 71.3 89.6 77.6 85.6
Table 5: Cross-Diffusion-Sources Evaluation on the GenImage Dataset Zhu et al. (2024).
Method
Mean Acc. of
16 sub-testsets
CNNDetectionWang et al. (2020) 70.8
LGradTan et al. (2023) 75.3
OjhaOjha et al. (2023) 78.4
FreqNetTan et al. (2024a) 78.1
NPRTan et al. (2024b) 81.8
ours 84.7
Table 6: The mean accuracy of all 16 generation models on GenImage Zhu et al. (2024) dataset.
DWT Window Tiling FFT WAttention mean Acc. mean A.P.
$\surd$ 67.2 70.8
$\surd$ $\surd$ 49.5 52.3
$\surd$ $\surd$ 82.3 87.6
$\surd$ $\surd$ $\surd$ 83.9 90.2
$\surd$ $\surd$ $\surd$ $\surd$ 84.7 90.9
Table 7: Ablation study on the GenImage Zhu et al. (2024).
Conclusion. To comprehensively evaluate the superiority of our method, we test it on the entire GenImage dataset Zhu et al. (2024) and calculated the average accuracy. The GenImage dataset offers extensive coverage, including content generated by GANs as well as high-definition images from diffusion models. As shown in Table 6, our method achieves optimal detection performance, with an accuracy of 84.7%, significantly surpassing the accuracy of NPR and FreqNet, which are 81.8% and 78.1%, respectively.
Figure 4: Logit Distributions of Extracted Forgery Features. We compare the state of the art NRP Tan et al. (2024b) and our method, both tuned with 4-class ProGAN Karras (2017) data. A total of four testing GANs and diffusion models are considered, including ProGAN Karras (2017), StyleGAN Karras (2019), IDDPM Nichol and Dhariwal (2021) and VQDM Gu et al. (2022), each randomly sampled 1k real and 1k fake images.
4.4 Qualitative Analysis
To further assess the generalization ability of our method, we visualize the logit distributions of NPR and our approach, as shown in Figure 4. This visualization highlights how effectively each trained model distinguishes between real and fake images, showcasing our method’s ability to generalize across diverse fake representations. From the visualization, it is evident that NPR struggles with unseen GAN or diffusion models, showing significant overlap between the logits of real and fake categories, often misclassifying fake images as ”real.” In contrast, our method demonstrates superior discrimination, effectively separating ”real” and ”fake” categories, even when encountering unseen sources.
4.5 Ablation Study
To thoroughly evaluate the effectiveness of each component in our method, we conduct ablation experiments by systematically removing some of the proposed modules. The results of these experiments, shown in Table 7, demonstrate the contribution of each module to deepfake detection. Notably, when we remove our main module (LoSFB), the performance of our method drops significantly, falling below 50%. This highlights the crucial role of local spatial-frequency domain feature extraction in detecting forged information.To further validate the effectiveness and rationale behind using the phase component of FFT to extract global information and employing sliding window tiling for local feature extraction in DWT, we remove these modules individually. As shown in the table, the accuracy decreased by 2.4% and 0.8%, respectively. These results emphasize the importance of both effectively extracting local texture information and subtle forged textures embedded in the global information.
5 Conclusion
In this paper, we propose a novel and effective method for deepfake detection, leveraging both local spatial-frequency information and global frequency domain features through two core modules: LoSFB and GloFB. This plug-and-play feature extractor is meticulously designed to capture a broader and more effective set of artifact features. Extensive experiments on 34 diverse generative models strongly demonstrate the generalization capability of our method.
6 Acknowledgments
This work is supported by the National Key R&D Program of China under grant 2021YFB2700900; by the National Natural Science Foundation of China under grant U22B2062, 62172232; by the Jiangsu Basic Research Programs-Natural Science Foundation under grant numbers BK20200039; by the Collaborative Innovation Center of Atmospheric Environment and Equipment Technology (CICAEET) fund.

### v2

@@
-Generalizable Deepfake Detection via Effective Local-Global Feature Extraction
+Dual Frequency Branch Framework with Reconstructed Sliding Windows Attention for AI-Generated Image Detection
 Abstract
-The rapid advancement of GANs and diffusion models has led to the generation of increasingly realistic fake images, posing significant hidden dangers and threats to society.
-Consequently, deepfake detection has become a pressing issue in today’s world.
-While some existing methods focus on forgery features from either a local or global perspective, they often overlook the complementary nature of these features.
-Other approaches attempt to incorporate both local and global features but rely on simplistic strategies, such as cropping, which fail to capture the intricate relationships between local features.
-To address these limitations, we propose a novel method that effectively combines local spatial-frequency domain features with global frequency domain information, capturing detailed and holistic forgery traces.
-Specifically, our method uses Discrete Wavelet Transform (DWT) and sliding windows to tile forged features and leverages attention mechanisms to extract local spatial-frequency domain information.
-Simultaneously, the phase component of the Fast Fourier Transform (FFT) is integrated with attention mechanisms to extract global frequency domain information, complementing the local features and ensuring the integrity of forgery detection.
-Comprehensive evaluations on open-world datasets generated by 34 distinct generative models demonstrate a significant improvement of 2.9% over existing state-of-the-art methods.
-1 Introduction
+The rapid advancement of Generative Adversarial Networks (GANs) and diffusion models has enabled the creation of highly realistic synthetic images, presenting significant societal risks, such as misinformation and deception.
+As a result, detecting AI-generated images has emerged as a critical challenge.
+Existing researches emphasize extracting fine-grained features to enhance detector generalization, yet they often lack consideration for the importance and interdependencies of internal elements within local regions and are limited to a single frequency domain, hindering the capture of general forgery traces.
+To overcome the aforementioned limitations, we first utilize a sliding window to restrict the attention mechanism to a local window, and reconstruct the features within the window to model the relationships between neighboring internal elements within the local region.
+Then, we design a dual frequency domain branch framework consisting of four frequency domain subbands of DWT and the phase part of FFT to enrich the extraction of local forgery features from different perspectives.
+Through feature enrichment of dual frequency domain branches and fine-grained feature extraction of reconstruction sliding window attention, our method achieves superior generalization detection capabilities on both GAN and diffusion model-based generative images.
+Evaluated on diverse datasets comprising images from 65 distinct generative models, our approach achieves a 2.13% improvement in detection accuracy over state-of-the-art methods.
+Index Terms:
+Deepfake detection, AI-generated image detection, AI security, Window attention.
+I Introduction
 With the rapid advancement of deep learning, generative models such as GANs and diffusion models have revolutionized content creation, finding widespread applications across industries including entertainment and media.
@@
 The increasing sophistication of deepfake technology has allowed the generation of highly realistic fake images and videos, making them increasingly difficult to detect and posing serious risks to societal trust, individual privacy, and security.
-In response, various deepfake detection mechanisms have been conceived.
-Early methods primarily focused on analyzing image features, such as texture, color, and lighting Yu et al.
-(2020); Liu et al.
-(2020), to identify forgery traces.
-However, these approaches were often limited to training and testing within a single domain, overlooking the critical issue of cross-domain generalization.
-In real-world scenarios, it is essential for detection models to effectively generalize to unseen deepfake sources, including those involving unfamiliar generative models or novel categories.
-Figure 1: Comparison with Previous Methods.
-Previous methods often extract only local information or global information to detect artifact information.
-However, these methods are not enough to fully capture effective artifact information.
-Recent methods also combine local and global information to detect forged images, but they only extract local information through simple cropping, destroying the relationship between local information.
-Compared with them, our method uses DWT and sliding window tiling to reasonably and effectively extract local time-frequency information, and combines the phase part of FFT to extract global frequency domain information.
-In recent years, substantial efforts have been devoted to forgery detection, particularly to identifying deepfake sources that are challenging to detect in real-world scenarios, as illustrated in Figure 1 (a).
-To enhance generalization performance, some studies emphasize fine-grained features, such as local artifacts Haliassos et al.
-(2021); Chen et al.
-(2022) and the relationships between pixel points Chen et al.
-(2021); Tan et al.
-(2024b).
-Others focus on global artifact traces, including global spatial information Tan et al.
-(2023); Ojha et al.
-(2023) and frequency-domain features Tan et al.
-(2024a).
-However, prioritizing either local or global information in isolation may result in the loss of complementary details, thereby limiting generalization capabilities.
-To address this limitation, some methods integrate both local and global features.
-For example, Chen and Yang (2020) extract local features through clipping in the spatial domain while leveraging global features in the frequency space.
-Similarly, Wu et al.
-(2024) employ a local attention module and facilitates interaction with features from distant adjacent PPG graphs to capture local and global information, respectively.
-However, these approaches often rely on coarse image block segmentation for local feature extraction, which can lead to forced separation and loss of relevant feature relationships.
-Accordingly, we propose a simple yet efficient feature extractor to capture global frequency domain and local spatial-frequency domain information, as illustrated in Figure 1 (b), which is plug-and-play and can be integrated with any CNN classifier.
-Specifically, our method introduces two key modules, DWT-Based Local Spatial-Frequency Domain Feature Capture Block (LoSFB) and FFT-Based Global Frequency Domain Feature Capture Block (GloFB), both using the Window Attention Block as the feature extractor.
-After performing the DWT, we concatenate features from four different frequency bands, each with dimensions $h\times w$.
-Then a $2\times 2$ sliding window is applied to extract the tiled features from each frequency band.
-Finally, the features within the window are tiled again, resulting in new features with dimensions $4\times(h\times w)$.
-At this time, when we extract the $4\times 4$ window features, they not only obtain features in four frequency domains, but also retain the local spatial-frequency information within the window, greatly improving the effect of artifact detection.
-Additionally, we incorporate the phase information from the FFT as a complement to the DWT, enabling the extraction of global frequency domain information.
-This helps guide the network to effectively learn global artifact patterns and identify more comprehensive traces of forgery.
-To fully assess the generalizability of our method, we evaluate it on large image datasets generated by 34 distinct models [^1] ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, SAN, AttGAN, BEGAN, CramerGAN, InfoMaxGAN, MMDGAN, RelGAN, S3GAN, SNGAN, STGAN, ADM, DDPM, IDDPM, LDM, PNDM, VQDiffusion, Stable-Diffusion v1, Stable-Diffusion v2, DALLE, Glide, Midjourney, Stable-Diffusion v1.4, Stable-Diffusion v1.5, Wukong, DALLE-2.
-Comprehensive experiments demonstrate that our method generalizes well across diverse and unseen sources, demonstrating a significant improvement over state-of-the-art methods by 2.9%.
-Our paper makes the following contributions:
-  • We propose a plug-and-play feature extractor that seamlessly integrates local spatial-frequency and global frequency-domain information.
-This method can be effortlessly incorporated into any CNN classifier and effectively improves the model’s generalization capability.
-  • We use sliding window interpretability to extract local spatial frequency information, and creatively use only phase spectrum to obtain global features to capture broader forgery artifacts.
-  • Our experiments validate the effectiveness of the proposed method, which exhibits strong generalization across image detection tasks involving 34 different generation models.
-The method achieves a notable 2.9% improvement over state-of-the-art approaches, highlighting its superiority and robustness.
-2 Related Work
-In this section, we present a concise review of deepfake detection methods, mainly dividing them into three main groups: deepfake detection based on local information, deepfake detection based on global information, and deepfake detection based on local and global information fusion.
-2.1 Deepfake Detection Based on Local Information
-Early researches on deepfake detection often focus on specific facial regions, such as the eyes and lips, to identify forgery traces Haliassos et al.
-(2021); Chen et al.
-(2022).
-However, as forgery techniques have advanced, these simple traces have been increasingly refined and obscured, rendering biometric-based detection methods insufficient for identifying sophisticated forged images.
-Experiments have shown that forgery traces are often embedded in fine-grained details, prompting a shift in focus toward extracting local artifact information.
-For instance, Chai et al.
-(2020) use limited receptive fields to capture local features, while Zhao et al.
-(2021) employ multiple spatial attention heads to enhance the network’s focus on distinct local regions.
-Other studies analyze forgery traces by examining relationships between local pixels.
-Chen et al.
-(2021) propose learning the similarity of local regions for a fine-grained assessment of patch similarity across scales, and Tan et al.
-(2024b) improve forgery detection generalization by identifying upsampling artifacts in nearby pixel grids.
-While capturing local artifact details, these methods overlook the fact that the overall image may still contain forged features, which are essential for generalization.
-2.2 Deepfake Detection Based on Global Information
-Early studies leveraged spatial information from generated images to detect fake content.
-For example, Yu et al.
-(2020); Liu et al.
-(2020) distinguish real and fake images using global information, such as color space and global texture.
-Some studies Yu et al.
-(2019) use GANs to generate global fingerprint information that may assist in detection.
-Additionally, Tan et al.
-(2023) and Ojha et al.
-(2023) extract artifact representations using global feature maps and global feature gradients, respectively.
-Moreover, since GAN-generated images often rely on upsampling operations that introduce artifacts at the frequency domain level, several studies have explored forgery detection in the frequency domain.
-For instance, Tan et al.
-(2024a) use FFT to extract global high-frequency information, uncovering more refined forgery features.
-While these methods capture sufficient global artifact information, they overlook the fact that effective artifact features often reside in subtle, localized regions.
-The neglect of local artifact information can significantly impair the generalization capabilities of forgery detection.
-2.3 Deepfake Detection Based on Local and Global Information Fusion
-Whether relying on global or local information alone, these methods detect forgery using only a single category of features, overlooking the significant information gain that arises from combining complementary features at different levels.
-Chen and Yang (2020) capture local features through spatial domain cropping and global features in the frequency domain.
-Xue et al.
-(2023) employ separate branches for local area feature extraction and global feature extraction, using the latter to obtain residual features.
-Wu et al.
-(2024) introduce a local attention module that interacts with distant neighboring PPG graphs to gather both local and global information.
-While these studies incorporate both global and local information, their approach to processing local information is limited by artificial segmentation of the face into regions, which leads to the loss of important forged details.
-3 Main Backbone
+To prevent the abuse of generative methods, a series of AI-generated image detection mechanisms have been conceived.
+Early methods primarily focused on analyzing image features, such as texture, color, and lighting [1, 2] to identify forgery traces.
+However, these methods are usually limited to the domain consistent scenario, where the training and testing sets come from the same type of generative method, resulting in poor generalized detection performance in cross-generative method scenarios.
+With the continuous deepening of research, it is widely recognized that artifacts are often confined to subtle and localized areas.
+Inspired by this, extensive research has been conducted to extract generalizable and fine-grained forgery features.
+Specifically, Tan et al.
+[3] extracted more generalized artifact information from the generator’s perspective by analyzing the relationship between adjacent pixels.
+Other works [4, 5, 6] have further amplified the local artifacts by means of segmenting patches, thus achieving better detection performance.
+In addition, some studies [7, 8, 9] captured forgery traces from the perspective of frequency domain to obtain more generalized detection performance.
+Although extracting local features has been widely adopted to capture generalized forgery artifacts, there are still some deficiencies as follow.
+1) The existing works all lack consideration of the importance differences and interdependencies of internal elements of local regions.
+Although the common attention mechanism [10, 11, 12] can adaptively assign different weights to different local regions, they are not sensitive enough to the importance and interrelationship among the internal elements within the local regions.
+2) Most existing works are limited to extracting local artifact features in a single frequency domain, resulting in the extracted features being insufficient for precisely capturing the general forgery traces caused by diverse generation methods.
+To address above issues, it is urgent for the proposed method to tackle the following two difficulties.
+1) How to model the importance difference and complex dependencies of elements within local regions?
+we adopt a sliding window to limit the attention mechanism to a local window and reconstruct the features within the window, making the attention mechanism more adaptively focused on pixel-level details while retaining the complex dependencies between pixels.
+Specifically, we concatenate features from the different frequency bands in each frequency domain.
+Then the nearby features are tiled in each frequency band.
+Finally, a sliding window is applied to extract the tiled features from all frequency bands.
+In this way, the reconstructed window features not only obtain features from the different frequency bands, but also retain the implicit relationships between adjacent elements.
+2) How to extract more abundant and effective local features to further amplify the forgery traces?
+It is well known that the DWT domain can reflect local spatial frequency features containing high-frequency and low-frequency texture information [13, 14, 15].
+In this paper, we find that each element of the phase part in the FFT domain contains the dependencies and structural relationship between the overall elements.
+That means it can capture the longer-range dependencies between the global elements, thereby enriching the extracted fine-grained local features from another perspective.
+To verify the rationality of above finding, we randomly selected 1k real images and 1k fake images from the dataset GenImage, and cropped them to the same size.
+We randomly exchange the phase part of the real image and the fake image, and use the swapped images as the input of the detector.
+The average prediction results of the images predicted as fake are shown in Figure 1.
+The smaller the prediction result is, the more likely the image is to be classified as a real image.
+We find that the detector tends to judge images with fake phase as fake category rather than fake amplitude.
+This indicates that the phase part contains more artifact traces.
+Moreover, the amplitude part mainly contains properties such as color and brightness, which is difficult to reflect forgery traces, as described in [16].
+Therefore, we design a dual frequency branch framework consisting of four frequency bands in the DWT domain and the phase part in the FFT domain to extract richer local artifact features from multiple perspectives.
+Based on this, a reconstructed sliding window attention mechanism is designed to obtain more fine-grained and generalized local artifact features, thus improving the detection performance across various generation methods.
+To fully assess the generalizability of our method, we evaluate it on large image datasets generated by 65 distinct generation models [^1] AttGAN, BEGAN, CramerGAN, InfoMaxGAN, MMDGAN, RelGAN, S3GAN, SNGAN, STGAN, ADM, DDPM, IDDPM, LDM, PNDM, VQDiffusion, SDv1, SDv2, Guided, Glide_50_27, Glide_100_10, Glide_100_27, LDM_100, LDM_200, LDM_200_cfg, DALLE, ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, ADM, Glide, Midjourney, SDv1.4, SDv1.5, VQDM, Wukong, DALLE2, R3GAN, StyleGAN3, StyleGAN-XL, StyleSwim, BlendFace, E4S, FaceSwap, InSwap, SimSwap, FLUX1-dev, Midjourney V6, GLIDE, DALLE-3, Imagen3, SD3, SDXL, BLIP, Infinite-ID, InstantID, IP-adapter, PhotoMaker, SocialRF, CommunityAI.
+Comprehensive experimental results show that our method achieves a significant improvement over state-of-the-art methods C2P_CLIP [17] and AIDE [18] by 5.8% and 7.0%.
+The overall contributions of the proposed method can be summarized as follows.
+Figure 1: Average Prediction Results of the Exchange Phase.
+The smaller the predicted number, the more likely the model is to classify the image as real.
+The detector consistently judged the images with false phase as false images rather than the images with false amplitude.
+  • We creatively reconstructed the sliding window attention mechanism, which effectively constructs the features within the window while limiting the attention within the local window, so that the attention mechanism can more accurately measure the importance and interdependence between internal elements in the local area.
+  • We discovered that the phase part of FFT domain can reflect the artifact traces precisely in addition to the high and low frequency information of the DWT domain, and further designed the dual frequency branch framework to extract richer local forgery features from multiple perspectives.
+  • Comprehensive experimental results have verified that our method exhibits strong generalization ability across 65 different generative methods.
+Compared with the state-of-the-art methods, the detection accuracy is improved by 2.13%, which further highlights its superiority and universality.
+II Related Work
+With the continuous development of generation methods, including GAN[19, 20, 21, 22], diffusion [23, 24, 25, 26], customized generation [27, 28, 29, 30, 31] and so on, the generated images are becoming more and more realistic, and the naked eye can no longer distinguish the authenticity of the images.
+In order to solve the future brought about by the above problems, many researchers are committed to the detection of AI-generated images.
+In this section, we will briefly introduce the existing detection methods from two aspects.
+II-A Exploration of Fine-Grained Features in AI-Generated Image Detection
+Early researches on AI-generated image detection often focused on specific facial regions, such as the eyes and lips, to identify forgery traces [32, 33].
+However, as forgery techniques have advanced, these simple traces have been increasingly refined and obscured, rendering biometric-based detection methods insufficient for identifying sophisticated forgery images.
+Experiments show that forgery traces are often embedded in fine details, which motivates people to turn to extracting local artifact information, including spatial and frequency domains.
+In terms of spatial domains, some researches narrow the focus to local areas to extract artifact information, including limited receptive fields [34, 35], cropping the image into smaller patches [4, 36, 6, 5], etc.
+Specifically, Chai et al.
+[34] employ limited receptive fields to identify patches that render images detectable, Zheng et al.
+[6] divides the image into multiple patches and shuffles them, and trains a block-based convolutional network for feature extraction.
+In addition, LGrad [8] employs pretrained CNN models as transform functions to convert images into gradients and leverages these gradients to present universal artifacts.
+NPR [3] improve generalization by identifying upsampling artifacts, such as pixel grid distortions, in nearby regions.
+In the frequency domain, F3-Net [37] introduces frequency component division and the frequency statistical difference between real images and forgery images into face forgery detection.
+FreDect [38] observes significant artifacts in the frequency domain of GAN-generated images, attributed to the upsampling operation in GAN architectures.
+Luo et al.
+[39] utilize multiple high-frequency features of images to enhance generalization performance.
+FreqNet [9] utilize the Fast Fourier Transform (FFT) to extract global high-frequency information, revealing refined forgery features such as periodic noise or anomalies invisible in the spatial domain.
+AIDE [18] improved the generalization ability of forgery detection by extracting the highest and lowest fine-grained patches through DCT to capture low-level fine-grained artifact information.
+Figure 2: Architecture of Our Method for Generalizable AI-Generated Image Detection.
+Specifically, we first design a reconstructed sliding window attention mechanism, which reconstructs the features within the sliding window and limits the attention mechanism to the local window range, forcing the attention to extract fine-grained forgery features while modeling the importance and dependencies between internal elements in the local area.
+At the same time, we designed a dual frequency branch framework, which facilitates the model to extract richer artifact traces from multiple perspectives by using DWT-based window tiling and the phase part of FFT.
+The final extracted features are passed into a classifier.
+II-B Application of Attention in AI-generated Image Detection
+Previous studies [10, 11, 12] has proven that: The attention mechanism can not only be trained to focus on important areas of the image, but also capture the long-range dependencies of the image, which is very useful in AI-generated image detection.
+Many works have explored this: Zhao et al.[40] employ multiple spatial attention heads, allowing the model to attend to distinct regions—such as the eyes and mouth—simultaneously, enhancing its ability to spot inconsistencies.
+Wu et al.[41] introduce a local attention module interacting with distant structures (e.g., PPG graphs), though further details from the original study would clarify this mechanism.
+In recent years, the attention mechanism has been widely integrated into advanced models such as Vision Transformer (ViT) [12] and CLIP [42], which not only improves the accuracy of detection, but also brings new possibilities for dealing with increasingly complex forgery techniques.
+For example, Ojha et al.
+[43] directly utilizes image features from the CLIP model for linear classification, LASTED [44] proposes designing textual labels to supervise the CLIP vision model through image-text contrastive learning, further advancing the field of AIGC detection, Fatformer [15] introduces a forgery-aware adapter to discern and integrate local forgery traces based on CLIP.
+CLIPMoLE [45] adapts a combination of shared and separate LoRAs within an MoE-based structure in deeper ViT blocks.
+What’s more, C2P_CLIP [17] explored the reasons and potential of CLIP’s effectiveness in AI-generated image detection, and for the first time integrated category common cues into the text encoder to inject category-related concepts into the image encoder, thereby improving detection performance.
+Existing attention mechanisms, such as Vision Transformers (ViT) and CLIP, typically capture global weights and long-range dependencies to extract broad semantic information.
+However, they lack sensitivity to the significance and interrelationships of elements within local regions, hindering their ability to effectively detect fine-grained forgery traces.
+Therefore, we propose a reconstructed sliding window attention mechanism, which enhancing the feature extraction capacity of the window attention block to model the importance and dependencies of local forgery features.
+In addition, a dual frequency branch network is designed to enrich the features within the window from different angles, which promotes the extraction of more comprehensive local forgery features in the window.
+III The Proposed Method
+III-A Main Backbone
 The overall structure of our method is shown in Figure 2.
-Inspired by Mao et al.
-(2024), we propose an efficient local-global feature extraction framework designed to capture both global and local information.
-Specifically, the framework consists of two main branches: LoSFB and GloFB.
-Both branches utilize the Window Attention Block to extract the most relevant features, which are then fused and passed to the classifier for final classification.
-Notably, our approach is compatible with any classifier, allowing us to focus primarily on the feature extraction process.
-Figure 2: Architecture of Our Method for Generalizable Deepfake Detection.
-In order to capture more comprehensive and reasonable forgery features, we use LoSFB and GloFB to capture local and global information, respectively.
-Specifically, we use DWT and sliding window tiling to construct local spatial-frequency features, and only use the phase part of FFT to construct global frequency domain features.
-Both of them use Window Attention Block to extract features.
-In addition, in order to weaken the information weight of global features, we add hyperparameters $\lambda$ to the global features when merging the two types of features.
-The final extracted features are passed into a classifier.
-3.1 Window Attention Block
-The Window Attention Block (WAttention) consists of two main components: (i) the Local Attention Block (LoAttention) and (ii) the Global Window MLP Layer (GMLP).
+Specifically, we first propose a reconstructed sliding window attention mechanism, which reconstructs the features within the window while limiting the attention to the local window, forcing the attention to extract fine-grained forgery traces while modeling the importance and interdependence between internal elements within the local range.
+In addition, in order to enrich the extraction of local features from different perspectives, we design a dual frequency branch framework, using DWT-based window tiling and phase part of FFT before using window attention.
+In this way, the window feature not only obtains the features of different frequency bands, but also retains the implicit relationship between adjacent elements, while complementing the long-range dependency between features, which improves the overall generalization ability of the model.
+Figure 3: After DWT, sliding window tiling is used to tile the features of each channel into a shape of $4\times(h\times w)$.
+In the subsequent LoAttention module, $\times 4$ 4 features are extracted as shown in the blue box, which not only obtain features from the different frequency bands, but also retain the implicit relationships between adjacent elements.
+III-B Reconstructed Sliding Window Attention Block
+The Reconstructed Sliding Window Attention Block (RSWAttention) consists of two main components: (i) the Local Attention Block (LoAttention) and (ii) the Global Window MLP Layer (GMLP).
 Before feeding the features into LoAttention and GMLP, we preprocess them to enhance their representational quality.
-A $1\times 1$ convolution layer is applied to extract cross-channel features, following by a $3\times 3$ depthwise separable convolution to better capture the local features within each channel.
+A $\times 1$ 1 convolution layer is applied to extract cross-channel features, following by a $\times 3$ 3 depthwise separable convolution to better capture the local features within each channel.
 The resulting convolved features are denoted as $X_{in}\in{\mathbb{R}}^{B\times C\times H\times W}$.
 LoAttention.
-Unlike traditional attention mechanisms, which are typically designed to capture global information, we employ a sliding window approach that divides the features into windows of size $b\times b$.
+The attention mechanism obtains better results by training different weights of features.
+However, the weights of traditional attention mechanisms are derived from the features of the entire image, which contains a lot of semantic information.
+This weight design may cause the model to be disturbed when extracting fine-grained forgery traces.
+Thus we designed LoAttention module.
+Specifically, we employ a sliding window approach that divides the features into windows of size $b\times b$.
 By restricting the attention module to these local windows, the focus shifts to effectively extracting local features.
@@
 $A_{i}=\text{Attention}(Q_{i},K_{i},V_{i})=\text{SoftMax}(Q_{i}K_{i}^{\top})V_{i}.$ (1)
-The results of all window attention operations can be expressed as $A=\{A_{1},A_{2},\cdots,A_{S}\}\in{\mathbb{R}}^{S\times C\times N}$.
+The results of all sliding window attention operations can be expressed as $A=\{A_{1},A_{2},\cdots,A_{S}\}\in{\mathbb{R}}^{S\times C\times N}$.
 GMLP.
@@
 After combining the outputs from LoAttention and GMLP via a dot product, we use the inverse window operation to restore the feature to its original size $B\times C\times H\times W$, followed by a convolution to map the features back to the original feature dimensions.
-The final output of the WAttention can be expressed as:
+The final output of the RSWAttention can be expressed as:
 $\text{X}_{W}=\text{conv}(\sum_{i=1}^{s}\text{Attention}(Q_{i},K_{i},V_{i})\cdot\sum_{i=1}^{s}G(\text{Linear}(V_{i}))).$ (3)
-3.2 DWT-Based Local Spatial-Frequency Domain Feature Capture Block
-Previous methods often leverage the Fast Fourier Transform (FFT) or Discrete Cosine Transform (DCT) to capture relationships between adjacent nodes in the frequency domain.
-However, these approaches can lead to the loss of positional information, which is crucial for extracting detailed features.
-To overcome this limitation, we utilize the Discrete Wavelet Transform (DWT), which decomposes the image into the spatial-frequency domain, represented as:
-$\text{DWT}(I)=\{I_{LL},I_{LH},I_{HL},I_{HH}\},$ (4)
-where $I_{LL},I_{LH},I_{HL},I_{HH}\in{\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}}$ represent the low-frequency component of the input and the high-frequency components in the vertical, horizontal, and diagonal directions, respectively.
-To obtain more comprehensive information, we fuse the data from the four frequency domains, resulting in new input features $\hat{I}\in{\mathbb{R}^{4\times C\times\frac{H}{2}\times\frac{W}{2}}}$.
-As illustrated in Figure 3, to extract detailed features, we divide each channel’s features into $2\times 2$ windows, with each window capturing localized spatial-frequency information within its respective frequency band.
-Each window is treated as a small square and tiled into rows within each band.
-This arrangement ensures that features within the $2\times 2$ windows are preserved in a structured format.
-Then we tile the features in the window here by rows.
-At this time, the features from each frequency band are tiled into one row in a specific manner, such that the four adjacent feature values correspond to the $2\times 2$ window from the original image.
-Additionally, as shown in Figure 3, we integrate features from all frequency bands to generate new input features ${\hat{I}}_{\text{input}}\in{\mathbb{R}}^{C\times 4\times\frac{H\times W}{4}}$.
-To further enhance the network’s performance, we apply LayerNorm Lei Ba et al.
-(2016), which ensures that the input data remains within a stable range, reduces dependency on initialization, and improves the stability of model training.
-The feature ${\hat{I}}_{\text{input}}$ is then fed into WAttention with a sliding window size set to $b=4$.
-As depicted in Figure 3, within each $4\times 4$ window, the representation includes not only the four adjacent feature values from the original image but also the feature information across four distinct frequency bands.
-This configuration provides WAttention with comprehensive and fine-grained spatial-frequency domain information, enabling more detailed feature extraction.
-Following this process, we apply the IDWT to reconstruct the features back into spatial domain representations, denoted as ${\hat{I}_{\text{DWT\_output}}}$.
-3.3 FFT-Based Global Frequency Domain Feature Capture Block
-Although artifact information is often concentrated in finer details, forged features can also be embedded in the global information of an image, remaining imperceptible to the naked eye.
-While DWT and sliding window tiling are effective for extracting local features, we also use FFT capture global frequency domain information from a complementary perspective.
-In this module, FFT is employed to transform the image into the frequency domain at the pixel level, enabling the extraction of global frequency domain features.
-Inspired by Zhao et al.
-(2024), we recognize that amplitude primarily reflects properties such as color and brightness, while phase encodes finer details, including potential artifacts.
-Based on this insight, we focus on extracting the phase features as input to WAttention, setting the window size to $b=8$.
-The amplitude remains unmodified during this process.
-After training the phase features, we use IFFT to recombine them with the original amplitude, transforming the features back into the spatial domain.
-The resulting features are denoted as ${\hat{I}_{\text{FFT\_output}}}$.
-Finally, we set a hyperparameter $\lambda$ to minimize the impact of global information extraction. The final feature is:
-${\hat{X}_{\text{output}}}=(1-\lambda)*\hat{I}_{\text{DWT\_output}}+\lambda*\hat{I}_{\text{FFT\_output}}.$ (5)
-Figure 3: After DWT, sliding window tiling is used to tile the features of each channel into a shape of $4\times(h\times w)$.
-In the subsequent LoAttention module, $4\times 4$ features are extracted as shown in the blue box, which not only contains local spatial frequency information, but also contains all different frequency band features.
-4 Experiments
-To verify the excellent deepfake detection performance of our method, we followe the paradigms of Ojha et al. (2023) and Wang et al. (2020).
-4.1 Dataset
-Training set.
-To ensure a consistent basis for comparison, we use the training set of ForenSynthsWang et al.
-(2020).
-The training set consists of 20 different categories, each containing 18,000 synthetic images generated using ProGAN, and an equal number of real images from the LSUN dataset.
-Following the paradigms of Ojha et al.
-(2023), we adopt specific 4-class training settings, denoted as (car, cat, chair, horse).
-Testing set. In order to comprehensively compare the effectiveness of our proposed method, we select different GAN and diffusion datasets for testing, which contain a wide range of real-world images.
-  • ForenSynths: The test set comprises 10 subsets, including: ProGAN Karras (2017), StyleGAN Karras (2019), StyleGAN2 Karras et al.
-(2020), BigGAN Brock (2018), CycleGAN Zhu et al.
-(2017), StarGAN Choi et al.
-(2018), GauGAN Park et al.
-(2019), Deepfake Rossler et al.
-(2019), WFIR and SAN Dai et al.
-(2019).
-  • GANGen-Detection Tan et al.
-(2024a): To evaluate more realistic scenarios, we extend our evaluation using 9 additional GAN-generated images.
-There are 4K test images for each model, with equal numbers of real and fake images.
-  • DiffusionForensics Wang et al.
-(2023): To expand the testing scope, we adopt the diffusions dataset of DIRE Wang et al.
-(2023) for evaluation, including ADM Dhariwal and Nichol (2021), DDPM Ho et al.
-(2020), IDDPM Nichol and Dhariwal (2021), LDM Rombach et al.
-(2022), PNDM Liu et al.
-(2022), VQ-Diffusion Gu et al.
-(2022), Stable Diffusion v1 Rombach et al.
-(2022), Stable Diffusion v2 Rombach et al.
-(2022).
-The real images are sampled from LSUN Yu et al.
-(2015) and ImageNet Russakovsky et al.
-(2015) datasets.
-  • UniversalFakeDetect Ojha et al.
-(2023): This test set contains images generated from Glide Nichol et al.
-(2021), DALL-Emini Ramesh et al.
-(2021), LDM Rombach et al.
-(2022).
-It adopts images of LAION Schuhmann et al.
-(2021) and ImageNetRussakovsky et al.
-(2015) datasets as the real data.
-  • AIGCDetectBenchmark Zhu et al.
-(2024): This dataset mainly adopts diffusion model to generate images, and is generated using more realistic advanced methods, including Midjourney, SDv1.4, SDv1.5 Rombach et al.
-(2022), ADM Dhariwal and Nichol (2021), GLIDE Nichol et al.
-(2021), Wukong, VQDM Gu et al.
-(2022), which is a huge challenge for deepfake detection.
-4.2 Implementation Details
-We design a lightweight CNN network using convolutional layers and Resnet network as classifiers.
-The network is trained using the Adam optimizerKingma (2014) with a learning rate of $2\times 10^{-4}$.
-The batchsize is set to 128, and we train the network for 90 epochs.
-We also use a learning rate decay strategy that reduces the learning rate by twenty percent after every ten epochs.
-What’s more, the hyperparameter $\lambda$ is set to 0.2.
-Our approach is implemented using PyTorch on an NVIDIA RTX A6000 GPU.
-To evaluate the performance of the proposed method, we follow the evaluation metrics used in baselines Ojha et al.
-(2023); Tan et al.
-(2024a), which include average precision score (A.P.) and accuracy (Acc.).
-4.3 Quantitative analysis
-We compare with the previous methods: CNNDetection Wang et al.
-(2020), Patchfor Chai et al.
-(2020), F3Net Qian et al.
-(2020), FrePGAN Jeong et al.
-(2022), GANDetection Mandelli et al.
-(2022), LGrad Tan et al.
-(2023), Ojha Ojha et al.
-(2023), FreqNet Tan et al.
-(2024a) and NPR Tan et al.
-(2024b).
-We conduct two and three sets of experiments on GAN and diffusion respectively.
-GAN.
-As shown in Table 1 and 2, our detection on GAN has achieved good results.
-Table 1 presents a detailed comparison of detector performance on the test set of ForenSynths Wang et al.
-(2020), highlighting higher mean accuracy (Acc.) and comparable mean average precision (A.P.) metrics.
-For unavailable test results, we use ”-” as a placeholder and exclude these values from the average calculations.
-Our method achieves an average accuracy of 86.9%, which is 1% and 1.7% higher than the latest methods, NPR and FreqNet, underscoring its effectiveness in generalizable deepfake detection.
-Furthermore, to evaluate cross-GAN-source generalization, we conducted tests on the GANGen-Detection dataset Tan et al.
-(2024a), as summarized in Table 2.
-Our method achieves an impressive average accuracy of 96.5%, significantly outperforming the best-performing baselines, NPR Tan et al.
-(2024b) and FreqNet Tan et al.
-(2024a), which achieved 94.8% and 94.0%, respectively.
-Diffusion.
-To further evaluate the generalization ability of our method, we train on ProGAN and test on images generated by various diffusion models.
-The results are presented in Tables 3 and 4.
-Despite being trained exclusively on ProGAN-generated images, our method demonstrates strong generalization capabilities across diverse diffusion models.
-Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 98.5% and 99.9%, respectively.
-Compared to NPR and FreqNet, our method improves mean Acc.
-by 2.9% and 17.8%, respectively.
-Additionally, on the UniversalFakeDetect dataset Ojha et al.
-(2023), our method achieves an average accuracy of 95.0%, representing a notable improvement over existing approaches.
-With the rapid advancements in tools like Midjourney and diffusion models, increasingly realistic fake images are being generated, posing significant challenges for deepfake detection.
-To address this, we conduct additional evaluations on the AIGCDetectBenchmark dataset Zhu et al.
-(2024), excluding some subsets of data previously tested.
-The results, summarized in Table 5, highlight the robustness of our approach.
-Our method achieves a notable improvement in mean accuracy, outperforming NPR and FreqNet by 3.7% and 7.8%, respectively, and reaching an accuracy of 77.6%.
-Method ProGAN GauGAN StyleGAN2 BigGAN StyleGAN CycleGAN StarGAN Deepfake WFIR SAN Mean
-Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
-CNNDetection 91.4 99.4 63.9 92.2 76.4 97.5 52.9 73.3 63.8 91.4 72.7 88.6 63.8 92.2 51.7 62.3 91.7 96.3 48.7 59.5 67.7 85.3
-Patchfor 97.8 100.0 57.2 55.4 83.6 98.5 64.7 69.5 82.6 93.1 74.5 87.2 100.0 100.0 85.0 93.2 - - 75.3 76.3 80.1 85.9
-F3Net 99.4 100.0 58.1 56.1 88.0 99.8 65.3 69.9 92.6 99.7 76.4 84.3 100.0 100.0 63.5 78.8 - - 47.3 46.7 76.7 81.7
-FrePGAN 99.0 99.9 60.3 77.1 84.1 98.6 69.2 71.7 80.7 89.6 71.1 74.4 99.9 100.0 70.9 91.9 - - - - 79.3 87.9
-LGrad 99.9 100.0 72.4 79.3 96.0 99.9 82.9 90.7 94.8 99.9 85.3 94.0 99.6 100.0 58.0 67.9 60.5 65.1 50.0 51.4 79.9 84.8
-Ojha 100.0 100.0 99.7 100.0 75.7 97.8 95.1 99.3 84.4 97.1 98.7 99.8 99.9 99.4 67.4 82.0 64.2 68.7 56.5 82.2 84.2 92.6
-FreqNet 99.6 100.0 93.4 98.6 88.0 99.5 90.5 96.0 90.2 99.7 95.8 99.6 85.7 99.8 88.9 94.4 48.0 49.6 71.9 80.1 85.2 91.4
-NPR 99.9 100.0 80.9 83.0 99.6 100.0 84.0 85.6 98.1 99.8 95.2 98.1 99.8 100.0 77.2 76.0 60.6 62.9 64.4 66.0 85.9 87.1
-ours 99.9 100.0 86.3 97.4 99.9 100.0 87.1 93.6 99.9 100.0 97.1 98.9 90.8 95.9 64.5 72.2 77.8 87.9 65.6 70.6 86.9 91.7
-Table 1: Cross-GAN-Sources Evaluation on the test set of ForenSynths Wang et al. (2020). Bold and underline represent the best and second-best performance, respectively.
+TABLE I: Cross-model Accuracy (Acc.) & Average Precision Score (A.P.) Performance on the GANGen-Detection Dataset [9].
 Method BEGAN SNGAN CramerGAN MMDGAN RelGAN STGAN S3GAN AttGAN InfoMaxGAN Mean
@@
 LGrad 69.9 89.2 78.0 87.4 50.3 50.3 57.5 67.3 89.1 99.1 54.8 68.0 78.5 86.0 68.6 93.8 71.1 82.0 68.6 80.8
-Ojha 88.9 96.3 87.7 96.5 89.8 99.1 89.7 99.0 94.0 98.1 82.6 91.5 94.7 98.9 90.0 96.8 87.4 96.5 89.4 97.0
+CLIPDetection 88.9 96.3 87.7 96.5 89.8 99.1 89.7 99.0 94.0 98.1 82.6 91.5 94.7 98.9 90.0 96.8 87.4 96.5 89.4 97.0
 FreqNet 98.8 100.0 85.4 90.4 95.2 98.2 95.2 98.2 100.0 100.0 98.8 100.0 88.3 94.3 89.8 98.8 94.0 97.5 94.0 97.5
 NPR 99.6 99.9 93.3 97.3 98.5 98.3 98.5 98.3 99.8 100.0 99.6 100.0 79.8 78.9 92.5 98.6 91.5 97.2 94.8 96.5
-ours 97.7 99.8 96.4 99.6 98.2 99.8 98.8 99.9 98.8 88.8 99.4 100.0 83.4 90.2 98.7 99.8 96.8 99.7 96.5 98.7
-Table 2: Cross-GAN-Sources Evaluation on the 9 GANs dataset Tan et al. (2024a).
+FatFormer 99.9 100.0 98.3 99.9 98.4 100.0 98.4 100.0 99.5 100.0 98.8 99.8 99.0 100.0 99.3 100.0 98.4 100.0 98.9 100.0
+Ladeda 100.0 100.0 99.0 99.9 97.9 99.9 99.0 99.9 100.0 100.0 99.8 100.0 80.5 84.9 100.0 100.0 98.4 99.9 97.2 98.3
+C2P_CLIP 94.9 100.0 98.4 99.9 98.4 100.0 98.4 100.0 92.0 99.8 97.6 99.6 99.0 100.0 90.4 99.8 98.4 100.0 96.4 99.9
+AIDE 57.9 79.5 76.1 79.2 82.0 88.1 82.2 88.3 95.2 96.5 92.7 98.1 69.0 78.8 96.6 98.4 80.7 82.8 81.4 87.8
+Ours 98.6 99.7 96.8 99.6 98.0 99.7 98.7 99.8 99.0 99.9 99.1 100.0 88.2 91.1 99.0 99.8 97.2 99.7 97.3 98.8
+TABLE II: Cross-model Accuracy (Acc.) & Average Precision Score (A.P.) Performance on the DiffusionForensics Dataset[46].
 Method DDPM VQ-Diffusion SDv1 ADM SDv2 IDDPM PNDM LDM Mean
@@
 LGrad 99.9 100.0 96.2 100.0 90.4 99.4 86.4 97.5 97.1 100.0 66.1 92.8 69.5 98.5 99.7 100.0 88.2 98.5
-Ojha 72.9 78.8 77.7 99.2 59.0 93.0 88.7 98.6 53.3 87.6 73.4 97.1 86.2 99.2 50.7 87.9 70.2 92.5
+CLIPDetection 72.9 78.8 77.7 99.2 59.0 93.0 88.7 98.6 53.3 87.6 73.4 97.1 86.2 99.2 50.7 87.9 70.2 92.5
 FreqNet 91.4 99.8 100.0 100.0 63.9 98.1 67.2 91.3 81.8 98.4 59.0 97.3 85.2 99.8 98.9 100.0 80.9 98.1
-NPR 99.4 100.0 100.0 100.0 97.3 99.8 88.7 98.6 96.4 100.0 87.5 98.3 97.5 100.0 100.0 100.0 95.8 99.6
-ours 99.4 100.0 99.9 100.0 95.6 99.7 96.3 99.4 99.2 100.0 99.4 100.0 99.4 100.0 99.8 100.0 98.7 99.9
-Table 3: Cross-Diffusion-Sources Evaluation on the test of DiffusionForensics Wang et al. (2023).
+NPR 99.4 100.0 100.0 100.0 97.1 99.7 88.4 97.9 96.4 100.0 87.5 98.0 97.5 100.0 100.0 100.0 95.8 99.4
+FatFormer 67.2 73.5 100.0 100.0 61.7 96.8 70.8 93.4 84.4 98.2 69.3 94.3 99.3 100.0 97.3 100.0 81.2 94.4
+Ladeda 98.4 99.9 100.0 100.0 97.1 99.8 83.9 96.9 94.1 100.0 93.9 98.8 93.7 99.9 100.0 100.0 95.1 99.4
+C2P_CLIP 73.4 76.2 95.8 99.7 78.9 99.2 68.8 95.3 66.7 94.8 80.7 94.9 84.2 97.2 97.3 99.7 80.7 94.6
+AIDE 86.5 92.5 92.1 99.4 91.5 98.6 85.7 94.1 88.6 95.1 87.8 94.8 89.6 97.8 87.4 95.0 88.6 95.9
+Ours 99.9 100.0 100.0 100.0 97.5 99.9 96.0 99.8 99.9 100.0 98.5 100.0 99.2 100.0 99.6 100.0 98.8 100.0
+III-C Dual Frequency Branch Feature Enrichment Module
+DWT-based window tiling branch.
+Previous methods often leverage the Fast Fourier Transform (FFT) or Discrete Cosine Transform (DCT) to capture relationships between adjacent nodes in the frequency domain.
+However, these approaches can lead to the loss of positional information, which is crucial for extracting detailed features.
+To overcome this limitation, we utilize the Discrete Wavelet Transform (DWT), which decomposes the image into the spatial-frequency domain, represented as:
+$\text{DWT}(I)=\{I_{LL},I_{LH},I_{HL},I_{HH}\},$ (4)
+where $I_{LL},I_{LH},I_{HL},I_{HH}\in{\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}}$ represent the low-frequency component of the input and the high-frequency components in the vertical, horizontal, and diagonal directions, respectively.
+To obtain more comprehensive information, we fuse the data from the four frequency domains, resulting in new input features $\hat{I}\in{\mathbb{R}^{4\times C\times\frac{H}{2}\times\frac{W}{2}}}$.
+As illustrated in Figure 3, to extract detailed features, we divide each channel’s features into $\times 2$ 2 windows, with each window capturing localized spatial-frequency information within its respective frequency band.
+Each window is treated as a small square and tiled into rows within each band.
+This arrangement ensures that features within the $\times 2$ 2 windows are preserved in a structured format.
+Then we tile the features in the window here by rows.
+At this time, the features from each frequency band are tiled into one row in a specific manner, such that the four adjacent feature values correspond to the $\times 2$ 2 window from the original image.
+Additionally, as shown in Figure 3, we integrate features from all frequency bands to generate new input features ${\hat{I}}_{\text{input}}\in{\mathbb{R}}^{C\times 4\times\frac{H\times W}{4}}$.
+To further enhance the network’s performance, we apply LayerNorm [47], which ensures that the input data remains within a stable range, reduces dependency on initialization, and improves the stability of model training.
+The feature ${\hat{I}}_{\text{input}}$ is then fed into RSWAttention with a sliding window size set to $b=4$.
+As depicted in Figure 3, within each $\times 4$ 4 window, the representation includes not only the four adjacent feature values from the original image but also the feature information across four distinct frequency bands.
+This configuration provides RSWAttention with comprehensive and fine-grained spatial-frequency domain information, enabling more detailed feature extraction.
+Following this process, we apply the IDWT to reconstruct the features back into spatial domain representations, denoted as ${\hat{I}_{\text{DWT\_output}}}$.
+FFT-based phase complement branch.
+Different from DWT, which captures local spatial frequency features, each element of the phase part in the FFT domain contains the dependencies and structural relationship between the overall elements, which helps the model capture the long-range dependencies between features.
+In addition, through our experimental analysis, we found that phase encodes finer details, including potential artifacts.
+Therefore, in order to enrich the extraction of local fine-grained features from another perspective, we only use the phase part of the FFT as another branch input of the RSWAttention, setting the window size to $b=8$.
+The amplitude remains unmodified during this process.
+After training the phase features, we use IFFT to recombine them with the original amplitude, transforming the features back into the spatial domain.
+The resulting features are denoted as ${\hat{I}_{\text{FFT\_output}}}$.
+Finally, we set a hyperparameter $\lambda$ to effectively combines the outputs of two branches. The final feature is:
+${\hat{X}_{\text{output}}}=(1-\lambda)*\hat{I}_{\text{DWT\_output}}+\lambda*\hat{I}_{\text{FFT\_output}}.$ (5)
+TABLE III: Cross-model Accuracy (Acc.) & Average Precision Score (A.P.) Performance on the UniversalFakeDetect Dataset [43].
 Method Guided Glide_50_27 Glide_100_10 Glide_100_27 LDM_100 LDM_200 LDM_200_cfg DALLE Mean
@@
 LGrad 71.8 76.0 91.6 96.0 90.9 95.5 88.5 94.2 95.7 99.3 94.7 99.1 95.1 99.0 89.9 97.8 89.8 94.6
-Ojha 69.7 87.6 79.1 99.8 77.9 94.5 78.5 95.3 95.0 99.3 99.0 99.3 74.0 92.5 87.3 97.5 82.6 95.7
+CLIPDetection 69.7 87.6 79.1 99.8 77.9 94.5 78.5 95.3 95.0 99.3 99.0 99.3 74.0 92.5 87.3 97.5 82.6 95.7
 FreqNet 67.3 75.7 86.7 96.3 87.9 96.4 84.5 96.0 97.9 99.9 97.5 99.9 97.4 99.9 97.4 99.8 89.6 95.5
-NPR 74.8 78.8 98.3 99.8 98.7 99.7 98.2 99.8 98.8 99.8 99.0 99.8 98.8 99.7 95.9 99.7 94.8 97.0
-ours 82.4 87.1 98.3 99.7 97.6 99.8 97.6 99.6 99.4 100.0 99.3 100.0 99.0 99.9 86.0 95.6 95.0 97.7
-Table 4: Cross-Diffusion-Sources Evaluation on the diffusion test set of Ojha Ojha et al. (2023).
+NPR 74.0 78.1 97.5 99.5 97.8 99.5 97.4 99.5 98.2 99.6 98.2 99.6 98.0 99.5 90.9 98.1 93.9 96.7
+FatFormer 76.1 92.0 94.7 99.4 94.2 99.2 94.4 99.1 98.7 99.9 98.6 99.8 94.9 99.1 98.8 99.8 93.8 98.5
+Ladeda 79.8 87.5 98.3 99.8 98.2 99.8 98.5 99.8 98.6 99.9 98.5 99.9 98.0 99.8 83.9 98.5 94.2 98.1
+C2P_CLIP 69.1 94.1 95.3 99.8 96.1 99.8 95.3 99.8 99.3 100.0 99.3 100.0 97.3 99.8 98.6 99.1 93.8 99.2
+AIDE 80.4 92.6 94.6 98.9 94.7 98.8 94.5 98.8 93.1 98.6 93.0 98.3 93.3 98.6 92.2 98.0 91.9 97.8
+Ours 95.2 97.8 98.0 99.9 97.7 99.9 97.2 99.7 99.4 100.0 99.5 100.0 98.9 99.9 85.2 97.1 96.4 99.3
+TABLE IV: Cross-model Accuracy (Acc.) & Average Precision Score (A.P.) Performance on the GenImage Datasets [48].
+Method ProGAN GauGAN StyleGAN2 BigGAN StyleGAN CycleGAN StarGAN Deepfake WFIR
+Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
+CNNDetection 97.9 99.9 66.1 93.5 67.0 97.1 55.9 80.5 70.1 97.0 76.5 91.5 72.9 93.2 51.7 66.5 62.1 89.4
+LGrad 99.9 100.0 72.4 79.3 96.0 99.9 82.9 90.7 94.8 99.9 85.3 94.0 99.6 100.0 58.0 67.9 60.5 65.1
+CLIPDetection 100.0 100.0 99.7 100.0 75.7 97.8 95.1 99.3 84.4 97.1 98.7 99.8 99.9 99.4 67.4 82.0 64.2 68.7
+FreqNet 99.6 100.0 93.4 98.6 88.0 99.5 90.5 96.0 90.2 99.7 95.8 99.6 85.7 99.8 88.9 94.4 48.0 49.6
+NPR 99.9 100.0 80.9 83.0 99.6 100.0 84.0 85.6 98.1 99.8 95.2 98.1 99.8 100.0 77.2 76.0 60.7 63.9
+FatFormer 99.9 100.0 99.4 100.0 98.8 99.9 99.5 99.9 97.2 99.8 99.3 100.0 99.8 100.0 93.2 98.0 88.2 98.5
+Ladeda 100.0 100.0 94.1 100.0 99.9 100.0 82.5 89.5 99.8 100.0 91.1 99.0 76.2 90.3 64.0 92.6 85.7 93.6
+C2P_CLIP 100.0 100.0 99.2 100.0 95.6 99.9 99.1 100.0 96.4 99.5 97.3 100.0 99.6 100.0 93.8 98.6 94.8 99.5
+AIDE 98.0 99.8 60.2 71.9 97.5 99.6 71.0 82.2 97.4 99.5 86.0 97.1 99.3 100.0 53.7 70.3 90.3 97.0
+Ours 99.8 100.0 85.1 96.7 99.9 100.0 87.1 95.1 99.8 100.0 97.9 98.8 94.9 99.2 64.2 71.8 52.2 46.8
 Method ADM Glide Midjourney SDv1.4 SDv1.5 VQDM Wukong DALLE2 Mean
 Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
+CNNDetection 52.1 67.8 51.7 67.1 51.7 64.7 50.6 59.2 50.6 59.2 51.5 63.5 50.5 58.5 50.1 49.2 61.3 76.7
 LGrad 65.8 70.2 60.6 73.8 64.6 70.3 65.9 68.7 66.3 69.1 67.2 70.6 63.2 65.0 66.9 86.9 65.1 71.8
-Ojha 67.6 87.2 61.7 85.5 56.7 74.5 63.1 87.2 62.8 86.3 84.9 96.8 71.2 90.9 49.5 60.2 73.5 83.6
+CLIPDetection 67.6 87.2 61.7 85.5 56.7 74.5 63.1 87.2 62.8 86.3 84.9 96.8 71.2 90.9 49.5 60.2 73.5 83.6
 FreqNet 83.3 91.5 81.7 88.9 69.9 79.0 64.3 74.3 64.9 75.5 81.7 89.6 57.7 67.0 55.1 54.6 69.8 77.6
-NPR 70.6 75.1 77.2 83.2 76.6 82.6 76.8 82.8 78.0 83.8 75.2 77.7 74.5 78.2 61.9 71.6 73.9 79.4
-ours 80.2 83.7 80.8 93.0 73.0 79.4 78.4 84.1 77.8 87.9 83.2 88.7 75.5 78.0 71.3 89.6 77.6 85.6
-Table 5: Cross-Diffusion-Sources Evaluation on the GenImage Dataset Zhu et al. (2024).
-Method
-Mean Acc. of
-16 sub-testsets
-CNNDetectionWang et al. (2020) 70.8
-LGradTan et al. (2023) 75.3
-OjhaOjha et al. (2023) 78.4
-FreqNetTan et al. (2024a) 78.1
-NPRTan et al. (2024b) 81.8
-ours 84.7
-Table 6: The mean accuracy of all 16 generation models on GenImage Zhu et al. (2024) dataset.
-DWT Window Tiling FFT WAttention mean Acc. mean A.P.
-$\surd$ 67.2 70.8
-$\surd$ $\surd$ 49.5 52.3
-$\surd$ $\surd$ 82.3 87.6
-$\surd$ $\surd$ $\surd$ 83.9 90.2
-$\surd$ $\surd$ $\surd$ $\surd$ 84.7 90.9
-Table 7: Ablation study on the GenImage Zhu et al. (2024).
-Conclusion.
-To comprehensively evaluate the superiority of our method, we test it on the entire GenImage dataset Zhu et al.
-(2024) and calculated the average accuracy.
-The GenImage dataset offers extensive coverage, including content generated by GANs as well as high-definition images from diffusion models.
-As shown in Table 6, our method achieves optimal detection performance, with an accuracy of 84.7%, significantly surpassing the accuracy of NPR and FreqNet, which are 81.8% and 78.1%, respectively.
+NPR 70.3 74.9 77.0 83.1 76.4 82.4 76.6 82.7 77.8 83.5 74.9 77.4 74.2 77.9 61.9 71.6 81.4 84.7
+FatFormer 78.5 91.8 88.1 96.0 56.1 62.8 67.8 81.1 68.1 81.1 86.9 97.0 73.1 85.9 69.7 81.9 86.1 92.6
+Ladeda 74.0 84.4 90.6 96.6 86.9 93.5 89.8 95.5 90.6 95.8 76.6 84.5 87.6 93.0 83.0 93.9 86.6 92.8
+C2P_CLIP 77.2 95.8 88.5 98.8 59.2 83.0 82.8 97.3 82.8 97.2 87.2 98.4 80.1 95.5 65.1 92.9 88.2 92.6
+AIDE 83.1 95.8 85.9 95.3 69.1 78.1 83.2 95.2 82.8 93.9 80.3 92.2 83.5 95.0 86.6 97.6 82.8 91.8
+Ours 94.2 97.3 95.2 98.7 94.7 97.8 94.2 97.6 94.5 98.0 96.5 98.3 93.5 96.9 96.8 99.2 90.6 93.6
+TABLE V: Cross-model Accuracy (Acc.) & Average Precision Score (A.P.) Performance on the AIGIBench Datasets [49].
+Method ProGAN R3GAN StyleGAN3 StyleGAN-XL StyleSwim WFIR BlendFace E4S FaceSwap
+Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
+CNNDetection 97.8 100.0 49.9 48.4 64.0 91.6 49.7 60.0 58.0 88.9 62.1 89.4 49.3 34.2 49.6 35.2 50.0 44.3
+LGrad 98.0 99.9 65.6 69.1 59.9 64.3 61.2 62.8 68.1 75.0 55.6 53.1 38.6 39.4 39.2 39.4 28.2 33.3
+CLIPDetection 99.5 100.0 93.0 98.6 68.4 79.0 87.1 95.0 94.5 99.3 91.0 97.8 47.2 42.7 72.4 82.5 69.3 79.4
+FreqNet 99.3 100.0 62.3 56.8 83.0 92.4 79.8 84.1 80.8 91.8 58.5 48.9 23.3 34.1 25.8 34.7 40.4 43.4
+NPR 99.8 100.0 80.8 75.6 84.1 89.4 74.5 75.9 84.7 91.7 77.8 77.0 33.9 33.3 34.0 34.4 41.1 39.5
+FatFormer 99.9 100.0 94.4 99.3 91.5 96.9 93.4 98.0 93.9 99.8 88.2 98.5 46.0 44.7 70.0 73.0 77.6 83.4
+LaDeDa 100.0 100.0 72.9 69.2 83.7 86.3 82.1 84.2 82.6 86.4 84.5 91.3 28.2 38.0 29.0 42.4 41.2 40.2
+C2P_CLIP 100.0 100.0 95.4 99.8 93.8 98.5 94.1 98.1 95.2 100.0 94.8 99.5 55.2 60.5 69.8 76.1 73.5 82.2
+AIDE 98.0 99.8 89.0 94.8 79.2 77.8 84.4 87.7 85.2 85.1 90.3 97.0 43.4 43.8 39.4 38.3 46.7 43.7
+Ours 99.8 100.0 94.6 94.9 93.2 98.2 82.3 86.4 94.7 98.6 52.2 46.8 44.1 44.0 45.1 46.4 50.8 66.8
+Method InSwap SimSwap FLUX1-dev Midjourney V6 GLIDE DALLE-3 Imagen3 SD3 SDXL
+Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
+CNNDetection 50.1 44.3 49.7 45.3 50.3 50.3 49.9 39.3 51.5 67.3 50.4 53.8 51.0 50.3 52.8 66.3 52.6 65.5
+LGrad 37.7 38.5 36.1 38.1 64.6 70.0 59.2 62.1 56.7 66.2 49.5 48.6 61.4 66.9 66.2 71.4 64.0 66.3
+CLIPDetection 48.9 48.2 50.1 51.5 46.3 39.6 44.8 35.6 69.4 81.7 46.8 38.4 48.1 44.1 49.0 48.6 50.7 53.6
+FreqNet 37.5 42.1 36.5 41.9 78.5 87.3 57.6 57.7 75.8 77.4 66.2 61.0 73.6 80.7 77.6 84.0 86.4 93.6
+NPR 41.0 36.9 40.5 38.5 87.2 93.6 69.7 71.1 78.2 85.2 57.6 61.9 84.6 91.0 86.4 93.5 87.0 93.7
+FatFormer 57.9 66.4 61.7 72.2 48.3 41.8 47.6 43.5 84.5 92.2 50.1 46.8 46.9 43.1 62.9 68.5 75.0 83.0
+LaDeDa 38.3 41.8 37.3 39.6 84.2 88.0 67.0 68.0 86.4 89.4 39.2 46.3 81.1 84.6 83.6 87.6 83.1 88.0
+C2P_CLIP 59.5 69.0 67.9 78.5 47.6 46.8 51.2 50.2 85.1 93.4 64.4 75.4 45.2 46.5 53.5 65.6 59.3 76.0
+AIDE 46.1 43.0 47.2 46.8 84.9 86.6 77.9 78.1 90.0 89.3 56.9 57.0 86.4 90.8 89.7 94.1 88.9 90.0
+Ours 52.1 63.7 48.4 57.7 93.2 96.8 86.4 92.7 94.2 98.4 46.2 49.0 93.6 96.0 95.4 97.6 95.9 98.7
+Method BLIP Infinite-ID InstantID IP-Adapter PhotoMaker SocialRF CommunityAI Mean
+Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
+CNNDetection 49.8 44.7 50.8 55.8 50.1 67.1 50.0 50.1 49.6 51.5 50.3 46.3 49.9 50.5 53.6 57.6
+LGrad 56.4 57.8 31.3 34.4 53.9 54.8 45.5 46.3 32.0 35.8 49.6 49.3 56.4 63.0 53.4 56.2
+CLIPDetection 57.8 68.9 72.4 84.2 67.4 79.5 62.3 73.9 47.0 41.4 49.2 46.2 49.9 50.0 63.3 66.4
+FreqNet 93.8 99.9 79.0 74.5 79.8 86.3 78.8 79.9 77.0 74.9 54.2 58.1 55.9 69.7 60.2 61.7
+NPR 94.0 97.0 62.0 62.8 71.3 75.9 79.0 81.1 46.4 52.0 54.5 58.4 54.4 60.0 68.2 70.8
+FatFormer 82.8 92.3 83.3 91.2 69.3 77.4 67.1 74.1 46.4 37.4 56.9 65.1 51.9 69.9 69.9 74.3
+LaDeDa 93.4 96.7 79.0 67.5 81.0 81.8 82.3 82.2 69.3 66.6 56.2 61.7 54.8 55.9 68.8 71.3
+C2P_CLIP 84.9 93.8 85.2 93.4 86.8 93.9 55.1 65.9 50.6 52.0 56.4 62.3 51.0 54.5 71.0 77.3
+AIDE 92.7 96.7 88.9 87.9 87.6 89.0 87.1 91.0 83.4 80.9 56.8 62.1 51.3 52.8 74.9 76.2
+Ours 99.2 99.7 90.0 91.0 94.6 97.7 93.3 96.6 92.7 95.2 60.8 70.8 54.6 66.6 77.9 82.0
+IV Experiments
+To verify the excellent AI-generated image detection performance of our method, we follow the paradigms of [43] and [50], which will be described in detail below.
+IV-A Dataset
+Training set.
+To ensure a consistent basis for comparison, we use the training set of ForenSynths[50].
+The training set consists of 20 different categories, each containing 18,000 synthetic images generated using ProGAN, and an equal number of real images from the LSUN dataset.
+Following the paradigms of [43], we adopt specific 4-class training settings, denoted as (car, cat, chair, horse).
+Testing set.
+In order to comprehensively compare the effectiveness of our proposed method, we select different GAN and diffusion model-based generated datasets for testing, which contain a wide range of real-world images.
+  • GANGen-Detection [9]: To evaluate more realistic scenarios, we extend our evaluation using 9 additional GAN-generated images.
+There are 4K test images for each model, with equal numbers of real and fake images.
+  • DiffusionForensics [46]: To expand the testing scope, we adopt the diffusions dataset of DIRE [46] for evaluation, including ADM [23], DDPM [51], IDDPM [52], LDM [24], PNDM [53], VQ-Diffusion [54], Stable Diffusion v1 [24], Stable Diffusion v2 [24].
+The real images are sampled from LSUN and ImageNet datasets.
+  • UniversalFakeDetect [43]: This test set contains images generated from Glide, DALLE-mini [55], LDM [24]. It adopts images of LAION and ImageNet datasets as the real data.
+  • AIGCDetectBenchmark [48]: This dataset comprehensively includes GAN and diffusion, and introduces some new diffusion generation methods as detection datasets, including ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, Midjourney, SDv1.4, SDv1.5, ADM, GLIDE, Wukong, VQDM, which is a huge challenge for deepfake detection.
+  • AIGIBench [49]: AIGIBench is the first benchmark to conduct a comprehensive evaluation of four critical components in the AIGI detection pipeline, and comprehensively simulates state-of-the-art image generation methods, including: (a) GAN-based noise-to-image generation (ProGAN, StyleGAN3, StyleGAN-XL, StyleSwim, R3GAN, and WFIR), (b) Diffusion for text-to-image generation (SD-XL, SD-3, DALLE-3, Midjourney-v6, FLUX.1-dev, Imagen-3, and GLIDE), (c) GANs for deepfake (BlendFace, E4S, FaceSwap, InSwap, and SimSwap), and (d) Diffusion for personalized generation (InstantID [27], Infinite-ID [28], PhotoMaker [29], BLIP-Diffusion [30], and IP-Adapter [31]).
+It also includes 2 general subsets featuring fake images collected from social media platforms, which called CommunityAI and SocialRF.
+IV-B Implementation Details
+We design a lightweight CNN network using convolutional layers and Resnet network as classifiers.
+The network is trained using the Adam optimizer[56] with a learning rate of $2\times 10^{-4}$.
+The batchsize is set to 128, and we train the network for 90 epochs.
+We also use a learning rate decay strategy that reduces the learning rate by twenty percent after every ten epochs.
+What’s more, the hyperparameter $\lambda$ is set to 0.4.
+For the loss function, we used BCELoss, which is commonly used in classification tasks.
+Our approach is implemented using PyTorch on an NVIDIA RTX A6000 GPU.
+To evaluate the performance of the proposed method, we follow the evaluation metrics used in baselines [43, 9], which include average precision score (A.P.) and accuracy (Acc.).
+IV-C Quantitative analysis
+We compare with the previous methods: CNNDetection [50], Patchfor [34], F3Net [37], FrePGAN [7], GANDetection [57], LGrad [8], CLIPDetection [43], FreqNet [9] ,NPR [3], FatFormer [15], Ladeda [35], C2P_CLIP [17], AIDE [18].
+We conduct comprehensive experiments on five datasets to verify the effectiveness of our method.
+Specifically, the results of Patchfor [34], F3Net [37], FrePGAN [7], GANDetection [57] all come from NPR [3], and we obtain results of FatFormer [15], C2P_CLIP [17] using the official pre-trained model.
+In addition, we retrain all other methods following the paradigms of [43], which adopts specific 4-class training settings, denoted as (car, cat, chair, horse).
+GANGen-Detection[9].
+Since our method uses ProGAN for training, we first perform cross-domain testing on GAN.
+As shown in Table I, our detection on GAN has achieved good results.
+Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 97.3% and 98.8%, respectively.
+Compared to Ladeda and C2P_CLIP, our method improves mean Acc.
+by 0.1% and 0.9%, respectively.
+However, on Fatformer, the accuracy is 1.1% higher than our method, which needs to be improved.
+DiffusionForensics [46] & UniversalFakeDetect [43].
+To further evaluate the generalization ability of our method, we train on ProGAN and test on images generated by various diffusion models.
+The results are presented in Tables II and III.
+Despite being trained exclusively on ProGAN-generated images, our method demonstrates strong generalization capabilities across diverse diffusion models.
+Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 98.8% and 100.0%, respectively.
+Compared with the best detection methods NPR and Ladeda, our method improves mean Acc.
+by 3.0% and 3.7%, respectively.
+However, methods Fatformer, C2P_CLIP, and AIDE, which have good results on other datasets, have an accuracy rate of less than 90% on this dataset, and even C2P_CLIP has only 80.7% accuracy.
+The most direct speculation for this result is that they used CLIP in the network skeleton, but the real reason is still worth further exploration.
+Additionally, on the UniversalFakeDetect dataset [43], our method achieves an average accuracy of 96.4%, 2.2% accuracy improvement over the best method.
+GenImage [48].
+Compared with testing only on diffusion or GAN, GenImage not only includes the classic GAN generation method, but also adds new and more realistic diffusion generation datasets, such as DALLE2.
+The test results are shown in Table IV.
+Our method achieves a notable improvement in mean accuracy, outperforming C2P_CLIP and Ladeda by 2.4% and 4.0%, respectively, and reaching an accuracy of 90.6%.
+AIGIBench [49].
+AIGIBench proposes a more challenging forgery detection dataset to evaluate the performance of forgery detection models from different perspectives.
+The test results are shown in Table V, our detection has achieved good results.
+Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 77.9% and 82.0%, respectively.
+Compared to Ladeda and C2P_CLIP, our method improves mean Acc.
+by 3.0% and 5.9%, respectively.
+Although our results are the best, there are still major problems and improvements: 1) Although our method has achieved an accuracy rate of more than 90% in most GAN and diffusion scenarios, it still has an accuracy rate of only about 60% in the dataset collected in unknown generated social scenarios.
+Other methods also have very low accuracy rates, which still needs further research and improvement.
+2) When it comes to face-swapping forgery detection, the performance is generally low, even less than 50%.
+Almost all methods have this problem, which is also a problem that needs to be considered and improved.
+TABLE VI: Module Ablation Studies on the Five Datasets.
+DWT Window Tiling FFT RSWAttention GANGen-Detection DiffusionForensics UniversalFakeDetect GenImage AIGIBench mean Acc. / A.P.
+49.6 / 49.6 48.6 / 48.6 47.9 / 47.9 49.4 / 49.4 49.6 / 49.6 49.0 / 49.0
+93.2 / 96.5 95.6 / 100.0 93.2 / 95.3 87.5 / 91.2 74.1 / 78.5 88.7 / 91.2
+95.2 / 97.1 96.9 / 100.0 95.0 / 97.4 88.7 / 92.5 75.5 / 79.8 90.3 / 93.4
+96.2 / 98.8 96.6 / 99.9 96.2 / 99.1 90.0 / 92.1 74.2 / 75.8 90.6 / 93.1
+96.6 / 98.8 98.8 / 100.0 96.4 / 99.3 90.6 / 93.6 77.9 / 82.0 92.1 / 94.1
+TABLE VII: RSWAttention Validity Studies on the Five Datasets.
+Attention GANGen-Detection DiffusionForensics UniversalFakeDetect GenImage AIGIBench mean Acc. / A.P.
+WAttention 70.7 / 82.9 65.8 / 89.5 73.2 / 88.4 71.5 / 81.2 58.3 / 63.2 67.9 / 81.0
+RSWAttention 96.6 / 98.8 98.8 / 100.0 96.4 / 99.3 90.6 / 93.6 77.9 / 82.0 92.1 / 94.1
+TABLE VIII: Phase-only Part of FFT Validity Studies on the Five Datasets.
+FFT GANGen-Detection DiffusionForensics UniversalFakeDetect GenImage AIGIBench mean Acc. / A.P.
+Phase Amp
+95.2 / 97.1 96.9 / 100.0 95.0 / 97.4 88.7 / 92.5 75.5 / 79.8 90.3 / 93.4
+96.6 / 98.8 98.8 / 100.0 96.4 / 99.3 90.6 / 93.6 77.9 / 82.0 92.1 / 94.1
+93.0 / 97.0 91.3 / 99.2 92.2 / 97.2 85.1 / 91.2 72.7 / 77.8 86.9 / 92.5
+TABLE IX: DWT Different Frequency Domain Subbands Studies on the Five Datasets.
+DWT GANGen-Detection DiffusionForensics UniversalFakeDetect GenImage AIGIBench mean Acc. / A.P.
+LL 65.7 / 79.9 63.6 / 84.7 68.5 / 82.2 66.0 / 78.0 57.3 / 62.3 64.2 / 77.4
+LH 75.9 / 89.3 68.3 / 87.0 74.3 / 89.1 70.4 / 82.3 60.2 / 66.0 69.8 / 82.7
+HL 78.9 / 88.3 69.7 / 86.2 74.5 / 85.6 71.8 / 82.7 61.7 / 66.2 71.3 / 81.8
+HH 77.3 / 88.7 67.2 / 87.2 74.6 / 85.9 71.9 / 84.6 61.5 / 67.7 70.5 / 82.8
+HL,HH 91.8 / 96.3 81.4 / 99.3 91.6 / 97.7 81.4 / 90.3 71.5 / 76.7 80.3 / 88.0
+LH,HL,HH 90.7 / 98.0 84.0 / 99.9 92.5 / 99.0 84.3 / 93.0 73.8 / 80.5 85.1 / 94.0
+LL,LH,HL,HH 96.6 / 98.8 98.8 / 100.0 96.4 / 99.3 90.6 / 93.6 77.9 / 82.0 92.1 / 94.1
+TABLE X: Hyperparameter $\lambda$ Ablation Experiment on the Five Datasets.
+$\lambda$ GANGen-Detection DiffusionForensics UniversalFakeDetect GenImage AIGIBench mean Acc. mean A.P.
+0.0 95.2 / 97.1 96.9 / 100.0 95.0 / 97.4 88.7 / 92.5 75.5 / 79.8 90.3 93.4
+0.1 95.7 / 98.9 97.2 / 100.0 95.8 / 99.0 89.4 / 93.0 77.0 / 81.1 91.0 94.4
+0.2 96.2 / 98.8 97.5 / 100.0 95.3 / 98.9 90.0 / 93.2 77.0 / 81.6 91.2 94.5
+0.4 96.6 / 98.8 98.8 / 100.0 96.4 / 99.3 90.6 / 93.6 77.9 / 82.0 92.1 94.7
+0.6 95.5 / 98.9 98.5 / 99.9 96.0 / 99.0 89.9 / 93.1 76.9 / 80.9 91.1 94.4
+0.8 95.0 / 97.8 93.4 / 100.0 93.7 / 98.5 89.9 / 94.2 76.5 / 80.2 89.7 94.1
+1.0 49.6 / 49.6 48.6 / 48.6 47.9 / 47.9 49.4 / 49.4 49.6 / 49.6 49.0 49.0
 Figure 4: Logit Distributions of Extracted Forgery Features.
-We compare the state of the art NRP Tan et al.
-(2024b) and our method, both tuned with 4-class ProGAN Karras (2017) data.
-A total of four testing GANs and diffusion models are considered, including ProGAN Karras (2017), StyleGAN Karras (2019), IDDPM Nichol and Dhariwal (2021) and VQDM Gu et al.
-(2022), each randomly sampled 1k real and 1k fake images.
-4.4 Qualitative Analysis
+We compare the state of the art NRP [3] and our method, both tuned with 4-class ProGAN [19] data.
+A total of four testing GANs and diffusion models are considered, including ProGAN [19], R3GAN [58], BLIP [30] and VQDM [54], each randomly sampled 1k real and 1k fake images.
+IV-D Qualitative Analysis
 To further assess the generalization ability of our method, we visualize the logit distributions of NPR and our approach, as shown in Figure 4.
@@
 From the visualization, it is evident that NPR struggles with unseen GAN or diffusion models, showing significant overlap between the logits of real and fake categories, often misclassifying fake images as ”real.” In contrast, our method demonstrates superior discrimination, effectively separating ”real” and ”fake” categories, even when encountering unseen sources.
-4.5 Ablation Study
-To thoroughly evaluate the effectiveness of each component in our method, we conduct ablation experiments by systematically removing some of the proposed modules.
-The results of these experiments, shown in Table 7, demonstrate the contribution of each module to deepfake detection.
-Notably, when we remove our main module (LoSFB), the performance of our method drops significantly, falling below 50%.
-This highlights the crucial role of local spatial-frequency domain feature extraction in detecting forged information.To further validate the effectiveness and rationale behind using the phase component of FFT to extract global information and employing sliding window tiling for local feature extraction in DWT, we remove these modules individually.
-As shown in the table, the accuracy decreased by 2.4% and 0.8%, respectively.
-These results emphasize the importance of both effectively extracting local texture information and subtle forged textures embedded in the global information.
-5 Conclusion
-In this paper, we propose a novel and effective method for deepfake detection, leveraging both local spatial-frequency information and global frequency domain features through two core modules: LoSFB and GloFB.
-This plug-and-play feature extractor is meticulously designed to capture a broader and more effective set of artifact features.
-Extensive experiments on 34 diverse generative models strongly demonstrate the generalization capability of our method.
-6 Acknowledgments
-This work is supported by the National Key R&D Program of China under grant 2021YFB2700900; by the National Natural Science Foundation of China under grant U22B2062, 62172232; by the Jiangsu Basic Research Programs-Natural Science Foundation under grant numbers BK20200039; by the Collaborative Innovation Center of Atmospheric Environment and Equipment Technology (CICAEET) fund.
+IV-E Module Ablation Studies.
+In order to comprehensively evaluate the effectiveness of our method, we conduct ablation studies on each module.
+The results of these experiments, shown in Table VI, demonstrate the contribution of each module to AI-generated image detection.
+Specifically, the experiment can prove that:
+FFT alone has negative impacts on Attention.
+When we remove our main module (DWT-Based Window Tiling), the performance of our method drops significantly, falling below 50%.
+In our design, FFT is integrated as a parallel branch to complement DWT, leveraging its phase information to capture long-range dependencies that enhance fine-grained feature richness for reconstructed sliding window attention.
+However, the localized focus of RSWAttention, optimized for fine-grained details, appears incompatible with FFT’s global frequency representation when not moderated by DWT’s multi-scale decomposition or Window Tiling’s spatial structuring.
+This misalignment suggests that FFT alone introduces overly coarse features, diluting the attention mechanism’s ability to discern subtle, localized artifacts critical for deepfake identification.
+Reasonable design of sliding window greatly improves the ability of attention extraction artifacts.
+Whether it is DWT or window tiling, the lack of any part will lead to a decrease in model accuracy.
+Without DWT, window attention cannot capture fine-grained artifacts.
+Removing Window Tiling yields a mean Acc.
+/ A.P.
+of 90.3% / 93.4%, a notable decline from the full model’s 92.1% / 94.1%, suggesting that Window Tiling preserves critical spatial relationships within local windows, allowing RSWAttention to exploit Dependencies between local features.
+Without it, the disruption of these relationships impairs the model’s ability to detect localized artifacts, which is very important in datasets with complex or subtle manipulations.
+Therefore, carefully designed in-window features based on DWT and window tiling are crucial for RSWAttention to extract fine-grained forgery features.
+It is effective to use the phase part of FFT as another branch.As shown in Table VI, when the phase part of FFT is excluded, the average accuracy of the model decreases 1.8%, indicating that FFT enhances the attention mechanism’s capacity to model interdependencies between distant image regions.
+FFT, as another frequency domain branch, enriches the local features within the window attention from different perspectives.
+IV-F Comparative Studies of Module Design
+To further verify that our proposed method helps the model extract more fine-grained multi-angle comprehensive forgery traces while modeling the importance and implicit dependencies between local features, we conducted further comparative studies on the modules, as follows.
+1) The reconstructed sliding window attention mechanism greatly improves the ability of attention to capture forgery features.
+As shown in Table VII, compared with WAttention, our RSWAttention improves the mean Acc.
+and mean A.P.
+by 24.2% and 13.1% respectively, which suggests that the RSWAttention greatly promotes the attention mechanism to capture fine-grained forgery traces within the local window while modeling the importance and interdependence between local features.
+2) Utilizing only the phase component of the FFT can reduce features not related to forgery compared to using the full FFT.
+As shown in Table VIII, if we use not only the phase part of the FFT but also the amplitude part as input to the network, the average accuracy of the model drops by 5.2%, confirming that amplitude introduces noise or irrelevant intensity information that confounds the detector.
+In contrast, the phase, encoding structural details like edges and shapes, aligns with the artifact-relevant features needed for detection.
+This finding corroborates prior image processing research, where phase information is prioritized over amplitude for structural analysis, reinforcing our design choice to leverage FFT’s phase as a complementary input branch.
+3) All subbands of DWT are indispensable.
+Each frequency domain subband of DWT has its own unique meaning, the low-frequency part retains the overall structure and semantic background of the image, and the high-frequency part captures subtle texture and edge information.
+Deleting any subband will reduce the comprehensiveness of the information in the window.
+As shown in Table IX, whether using only one frequency domain subband or omitting some frequency domain subbands, the performance of the model will be degraded.
+IV-G Hyperparameter $\lambda$ Ablation Study.
+In order to more comprehensively evaluate the effectiveness of our method, we also conduct ablation experiments on the hyperparameter $\lambda$, which can be shown in Table X.
+When our parameter $\lambda$ is set to 0.4 as set in the paper, the best results are achieved on all datasets.
+V Conclusion
+In this paper, we propose a novel and effective method for AI-generated image detection.
+Specifically, we designed a reconstructed sliding window attention mechanism, which reconstructs the features within the window and limits the attention to the local window, forcing the attention to extract fine-grained features while modeling the importance and dependencies between internal elements in the local area.
+In addition, we design a dual frequency branch framework consisting of four frequency bands in the DWT domain and the phase part in the FFT domain to extract richer local artifact features from multiple perspectives.
+Extensive experiments on 65 diverse generative models strongly demonstrate the generalization capability of our method.